### v1

# Each of those eight coalition logics is also determined by six other kinds of models

Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism. They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions they meet. In this paper, we show that each of the eight coalition logics is also determined by the following six kinds of models, with the respective properties: single-coalition-first action models; single-coalition-first neighborhood models; clear grand-coalition-first action models; clear single-coalition-first neighborhood models; tree-like grand-coalition-first action models; tree-like single-coalition-first neighborhood models.
Keywords: coalition logics, action models, neighborhood models, seriality, independence of agents, determinism
1 Introduction
1.1 Eight coalition logics determined by eight classes of general concurrent game models
Coalition Logic $\mathsf{CL}$ proposed by Pauly ([10, 11]) is a central logic in logical research on strategic reasoning.
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{C}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{C}$ ensures $\phi$.
$\mathsf{CL}$ has two kinds of models.
The first kind is concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents who form coalitions; for every coalition, there is an availability function, which specifies available joint actions of the coalition at states; for every coalition, there is an outcome function, which specifies possible outcome states of joint actions of the coalition.
The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action in the model.
The second kind is alpha neighborhood models. Roughly, in an alpha neighborhood model: there are some states; there are some agents who form coalitions; for every coalition, there is an alpha neighborhood function, which specifies the alpha powers of the coalition. An alpha power is a scope into which the coalition can force the next moment to fall.
The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
[9] proposed that concurrent game models have three too-strong assumptions: the independence of agents, seriality, and determinism. Based on general concurrent game models, which do not have the three assumptions, a Minimal Coalition Logic $\mathsf{MCL}$ was presented.
Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some special kinds of situations. Considering which of the three properties we want to keep, there are eight coalition logics in total. [8] showed the completeness of the eight logics in a uniform way. It should be pointed out that the completeness $\mathsf{CL}$ is already given in [11, 5] and the completeness of $\mathsf{MCL}$ is already given in [9].
General concurrent game models are called grand-coalition-first action models in this paper.
1.2 Our work
In this work, we show that each of the eight coalition logics is also determined by six other kinds of models with respective properties.
  • The first kind is single-coalition-first action models. In this kind of models, the outcome function for single coalitions determines the outcome functions for other coalitions and the availability functions for all coalitions.
    This kind of models makes good sense for situations where we know the space of future possibilities and how agents’ behavior changes the future beforehand.
  • The second kind is single-coalition-first neighborhood models. In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition. An actual power is a minimal scope in which the coalition can force the next moment to fall. In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
  • The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition do not have the same outcome.
  • The fourth kind is clear single-coalition-first neighborhood models. In this kind of models, the powers of an agent are pairwise disjoint.
  • The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history.
  • The sixth kind is tree-like single-coalition-first neighborhood models. In this kind of models, every state has a determined history, too.
The rest of the paper is structured as follows.
  • In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from the former to the latter.
  • In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models.
  • In Section 4, we discuss compositionality of availability functions and outcome functions in action models.
  • In Section 5: we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models, and show they can represent single-coalition-first action models.
  • In Section 6: we define clear grand-coalition-first action models, and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models, and show that they can represent clear grand-coalition-first action models.
  • In Section 7, we define tree-like grand-coalition-first action models, and show they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models, and show they can represent tree-like grand-coalition-first action models.
  • In Section 8, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties.
  • We make some concluding remarks in Section 9.
2 General settings of coalition logics
2.1 Language
Let $\mathtt{AP}$ be a countable set of atomic propositions and $\mathtt{AG}$ be a finite nonempty set of agents. The subsets of $\mathtt{AG}$ are called coalitions. $\mathtt{AG}$ is called the grand coalition.
Definition 1 (The language $\Phi$).
The language $\Phi$ is defined as follows, where $p$ ranges over $\mathtt{AP}$ and $\mathrm{C}\subseteq\mathtt{AG}$:
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates that some available joint action of $\mathrm{C}$ ensures $\phi$. The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual. Define the dual $[\langle\mathrm{C}\rangle]\phi$ of $\langle[\mathrm{C}]\rangle\phi$ as $\neg\langle[\mathrm{C}]\rangle\neg\phi$, indicating every available joint action of $\mathrm{C}$ enables $\phi$.
2.2 Action semantics
Let $\mathtt{AC}$ be a nonempty set of actions. For any $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is a set of joint actions of $\mathrm{C}$. For any $a\in\mathtt{AG}$, we write $\mathtt{JA}_{a}$ instead of $\mathtt{JA}_{\{a\}}$. We define $\mathtt{JA}=\bigcup\{\mathtt{JA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
Definition 2 (Action models).
An action model is a tuple $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},$$\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an available action function for $\mathrm{C}$.
  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Different logics might put different constraints on action models.
Definition 3 (Action semantics of $\Phi$).
$\mathtt{AM},s\Vdash\top$ always
$\mathtt{AM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{AM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{AM},s\Vdash\phi$
$\mathtt{AM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{AM},s\Vdash\phi$ and $\mathtt{AM},s\Vdash\psi$
$\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
It can be verified:
$\mathtt{AM},s\Vdash[\langle\mathrm{C}\rangle]\phi$Leftrightarrow for all $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s,\mathrm{C})$, there is $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ such that $\mathtt{AM},t\Vdash\phi$
2.3 Neighborhood semantics
Definition 4 (Neighborhood models and alpha neighborhood models).
A neighborhood model is a tuple $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $\mathrm{C}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
A neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is called an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is closed under supersets.
Different logics may put different constraints on neighborhood models.
Definition 5 (Neighborhood semantics of $\Phi$).
$\mathtt{NM},s\Vdash\top$ always
$\mathtt{NM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{NM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{NM},s\Vdash\phi$
$\mathtt{NM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{NM},s\Vdash\phi$ and $\mathtt{NM},s\Vdash\psi$
$\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
The following fact indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing which formulas it satisfies.
Fact 1.
Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\mathtt{AG},$$\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
Then, for every $s\in\mathtt{ST}$ and $\phi\in\Phi$, $\mathtt{NM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
2.4 Transformation of action semantics to neighborhood semantics
Definition 6 (Actual and alpha effectivity functions of action models).
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model, and $\mathrm{C}\subseteq\mathtt{AG}$.
The function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
$\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
The function $\mathtt{LE}_{\mathrm{C}}$ defined as follows is called the alpha effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
$\mathtt{LE}_{\mathrm{C}}(s)=\{\mathrm{Y}\subseteq\mathtt{ST}\mid\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathrm{Y}\text{ for some }\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
In a sense, actual effectivity functions and alpha effectivity functions of action models contain important information on coalition powers.
Actual effectivity functions are more transparent in the following sense: each element of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions.
Definition 7 (Action models $z$-representable by neighborhood models and $\alpha$-representable by alpha neighborhood models).
We say a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $z$-represents an action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$.
We say a class of neighborhood models $\mathbf{NM}$ $z$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
We say an alpha neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $\alpha$-represents an action model $\mathtt{AM}$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{LE}_{\mathrm{C}}$.
We say a class of alpha neighborhood models $\mathbf{NM}$ $\alpha$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}$, and (2) every alpha neighborhood model in $\mathbf{NM}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
Intuitively, that a neighborhood model $\mathtt{NM}$ represents an action model $\mathtt{AM}$ means that $\mathtt{NM}$ contains all the information of $\mathtt{AM}$ to evaluate formulas of $\Phi$.
Theorem 1 (Representability implies equivalence).
  1. (1)
    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
  2. (2)
    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
Proof.
Here, we show only (1). We can show (2) in a similar way.
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model such that $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$. Let $s\in\mathtt{ST}$ and $\phi\in\Phi$. We want to show $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
The proof is by induction of $\phi$.
It is trivial when $\phi$ is a proposition letter p, $\phi$ is $\top$, $\phi$ is $\neg\psi$ and $\phi$ is $\psi\wedge\theta$. We are done if we can show when $\phi$ is $\langle[\mathrm{C}]\rangle\psi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
$\Rightarrow$
Assume $\mathtt{AM},s\Vdash\phi$. Then, $\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$. Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$. Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$. Then, $\mathtt{NM},s\Vdash\phi$.
$\Leftarrow$
Assume $\mathtt{NM},s\Vdash\phi$. Then, $\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$. Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$. Then, $\mathrm{Y}\in\mathtt{AE}_{\mathrm{C}}(s)$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$. Then, $\mathtt{AM},s\Vdash\phi$.
3 Eight coalition logics determined by eight classes of grand-coalition-first action models
3.1 Eight classes of grand-coalition-first action models
Definition 8 (Grand-coalition-first action models).
A grand-coalition-first action model is a tuple $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathsf{out}_{\mathtt{AG}}:\mathtt{ST}\times\mathtt{JA}_{\mathtt{AG}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathtt{AG}$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 9 (Outcome functions and availability functions).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}.$
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
Outcome functions in grand-coalition-first action models are given in a from-whole-to-part way: the outcome states of a coalition $\mathrm{C}$ performing a joint action $\sigma_{\mathrm{C}}$ at a state $s$ are determined by the outcome states of the grand coalition performing action profiles containing $\sigma_{\mathrm{C}}$ at $s$.
Fact 2.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\mid\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ for some }\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(s)\}$.
This fact indicates that availability functions in grand-coalition-first action models also are given in a from-whole-to-part way: available joint actions of a coalition $\mathrm{C}$ at a state are determined by available joint actions the grand coalition at the state.
Definition 10 (Successor functions).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Define the successor function $\mathtt{suc}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}.$
For any set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$.
Fact 3.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathtt{suc}(s)$.
Proof.
$\Rightarrow$
Let $\mathrm{C}\subseteq\mathtt{AG}$. Let $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$. Then, there is $\sigma_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$. That is $w\in\mathtt{suc}(s)$.
$\Leftarrow$
Let $w\in\mathtt{suc}(s)$. Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$.
Definition 11 (Seriality, independence, and determinism of grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say:
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, ${\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})}$ is a singleton.
We use $\mathrm{ES}$ to indicate the set of the eight strings: $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$. We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively. We use the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ to correspond to the eight combinations of the three properties, respectively.
For every $\mathtt{X}\in\mathrm{ES}$ and grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$, we say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action $\mathtt{X}$-model if $\mathtt{G}\textbf{-}\mathtt{AM}$ has the properties corresponding to $\mathtt{X}$.
3.2 Eight axiomatic systems
Definition 12 (An axiomatic system for $\mathsf{MCL}$).
Axioms:
Tautologies ($\mathtt{A}\text{-}\mathtt{Tau}$): all propositional tautologies
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle[\mathrm{C}]\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{C}]\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
Inference rules:
Modus ponens ($\mathtt{MP}$): $\dfrac{\phi,\phi\rightarrow\psi}{\psi}$
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle[\mathrm{C}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi}$
We let the following formulas respectively correspond to $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$.
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$
where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
Definition 13 (Axiomatic systems for $\mathsf{MCL+X}$).
For all $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inferences of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [8]).
The axiomatic system for $\mathsf{MCL+X}$ given in Definition 13 is sound and complete with respect to the set of valid formulas in $\Phi$ relative to the class of grand-coalition-first action $\mathtt{X}$-models.
Remarks
The class of grand-coalition-first $\mathsf{SID}$-models is $\alpha$-representable ([11, 5]), and the class of grand-coalition-first $\mathsf{SD}$-models is $\alpha$-representable ([14]). Are the other six classes of grand-coalition-first action models $\alpha$-representable? Are the eight classes of grand-coalition-first action models $z$-representable? It is not known yet.
4 About compositionality of availability functions and outcome functions in grand-coalition-first action models
We say availability functions of a grand-coalition-first action model is compositional if the availability function of a coalition can be derived from the availability functions of its members in a natural way. In a similar sense, we talk about compositionality of outcome functions of grand-coalition-first action models.
4.1 Neither outcome functions nor availability functions in grand-coa-lition-first action models are compositional
The following two facts are easy to verify, and we skip their proofs.
Fact 4.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
Then:
  1. (1)
    1. (a)
      for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
    2. (b)
      for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})$.
  2. (2)
    1. (a)
      for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
    2. (b)
      for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, then $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
Note that none of the converses of these items holds. Figure 1 indicates a counter-example for the first two items and Figure 2 indicates a counter-example for the last two items.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{1}\atop\alpha_{2}\beta_{2}$$\alpha_{1}\beta_{2}\atop\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
Figure 1: An example of that outcome functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{2}$$\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
Figure 2: An example of that availability functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
4.2 Availability functions in grand-coalition-first action models meeting independence of agents are compositional
The following fact is easy to check:
Fact 5.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action $\mathtt{I}$-model.
Then:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$: $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$ if and only if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$: $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ if and only if $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
Remarks
Models of Coalition Logic are grand-coalition-first action $\mathtt{SID}$-models. In those models, availability functions are compositional, but outcome functions might not be. Figure 1 indicates a counter-example.
4.3 Outcome functions are compositional in many situations
There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. Outcome functions are compositional in these situations.
What follows are two examples, and the second is from [3].
Example 1.
Adam and Bob control a ship lock, which has two doors. Adam can close or open the front door and Bob can close or open the back door. They cannot open the two doors at the same time. This scenario can be represented by the grand-coalition-first action model depicted in Figure 3.
$s_{1}$$s_{2}$$s_{3}$$\{c_{f},c_{b}\}$$\{c_{f}\}$$\{c_{b}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{open_{b}})$$(\mathtt{skip},\mathtt{close_{b}})$$(\mathtt{open_{f}},\mathtt{skip})$$(\mathtt{close_{f}},\mathtt{skip})$
Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1, where outcome functions are compositional. Here $c_{f}$ and $c_{b}$, respectively, express the propositions the front door is closed and the back door is closed. Here are some instances of outcome functions: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{2},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{2},\mathtt{skip})=\{s_{2}\}\cap\{s_{1},s_{2}\}=\{s_{2}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
Example 2.
A system has two processes, $a$ and $b$. The process $a$ can assign values $1$ or 00 to the variable $x$, and the process $a$ can assign values $1$ or 00 to the variable $y$. When $x=0$, then $a$ can leave the value of $x$ unchanged or change it to $1$. When $x=1$, then $a$ leaves the value of $x$ unchanged. Similarly, when $y=0$, then $b$ can leave the value of $y$ unchanged or change it to $1$. When $y=1$, then $b$ leaves the value of $y$ unchanged. Figure 4 indicates a grand-coalition-first model for this scenario, where outcome functions are compositional.
$s_{1}$$s_{2}$$s_{3}$$s_{4}$$\{x_{1},y_{1}\}$$\{x_{1},y_{0}\}$$\{x_{0},y_{1}\}$$\{x_{0},y_{0}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{x:=1},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$
Figure 4: This figure indicates a grand-coalition-first action model where outcome functions are compositional. Here are some instances: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
The following settings about agency can be found in the literature, such as [6] and [7]: there are some agents; every atomic proposition is controlled by at most one agent. In these settings, outcome functions are compositional.
5 Single-coalition-first action models and single-coalition-first neighborhood models
5.1 Single-coalition-first action models
Definition 14 (Single-coalition-first action models).
A single-coalition-first action model is a tuple $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathsf{out}_{a}:\mathtt{ST}\times\mathtt{JA}_{a}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $a$ such that for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 15 (Outcome functions and availability functions).
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model.
  • For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
  • For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
Clearly, the outcome functions in single-coalition-first action models are compositional.
Let $\mathtt{AC}$ be a set of actions and $\mathtt{JA}_{\mathrm{C}}$ be the set of joint actions of $\mathrm{C}$. For any $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}$, if for any $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma^{\prime}_{\mathrm{C}}|_{a}$.
Theorem 3.
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model. The following three conditions are equivalent:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
    1. (a)
      for all $a\in\mathtt{AG}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$;
    2. (b)
      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$, where $\mathrm{C}\neq\emptyset$.
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$:
    1. (a)
      for all $a\in\mathtt{AG}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$;
    2. (b)
      $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
  3. (3)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
    1. (a)
      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$;
    2. (b)
      for all fusion $\sigma^{\prime\prime}_{\mathrm{C}}$ of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, for all $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$, then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$. It suffices to show (2b).
Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{\mathrm{D}}$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. By (1b), $t\in\mathsf{out}_{b}(s,\sigma_{b})$, where $b\in\mathrm{D}$. Then, $t\in\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}$. Note $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{D}=\emptyset$. Similarly, we can show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{C}\neq\emptyset$ and $\mathrm{D}\neq\emptyset$.
Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\mathrm{D}=\{b_{1},\dots,b_{m}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$ and $\sigma_{\mathrm{D}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$ and $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$. Then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$.
Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}\cup\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
(2) $\Rightarrow$ (1)
Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. It suffices to show (1b).
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$.
Note $\mathrm{C}=\{c_{1}\}\cup\mathrm{C}-\{c_{1}\}$. By (2b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ = $\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$.
Note $\mathrm{C}-\{c_{1}\}=\{c_{2}\}\cup\mathrm{C}-\{c_{1},c_{2}\}$. By (2b), $\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$ = $\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2}\}})$.
Finally, we get
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}}(s,$ $\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}})$.
That is, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap$ $\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$.
Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
(2) $\Rightarrow$ (3)
Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
First, we show (3a).
  • Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. We want to show $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. We are done if we can show there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. It suffices to show that there is a joint action $\sigma_{\overline{\mathrm{C}}}$ of $\overline{\mathrm{C}}$ such that $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
    Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\emptyset$. Then, $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$. Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, there are $\sigma_{a_{1}},\dots,\sigma_{a_{n}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}}),\dots,w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
    Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, $w\in\bigcup\{\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\mid\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}\}$. Then, $w\in\mathtt{suc}(s)$. Assume ${\overline{\mathrm{C}}}=\emptyset$. Then, $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathtt{suc}(s)$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$. Assume ${\overline{\mathrm{C}}}\neq\emptyset$. Assume ${\overline{\mathrm{C}}}=\{b_{1},\dots,b_{n}\}$. As for every $b\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathtt{suc}(s)$, there are $\sigma_{b_{1}},\dots,\sigma_{b_{n}}$ such that $w\in\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}}),\dots,w\in\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{n}}$. By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
  • Let $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. We want to show $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. By (2b), $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$. Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
Second, we show (3b). Let $\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, and $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{\mathrm{C}}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$. Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$, where $\sigma_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma_{c_{n}}\in\mathtt{JA}_{c_{n}}$, $\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime}_{c_{n}}$, where $\sigma^{\prime}_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma^{\prime}_{c_{n}}\in\mathtt{JA}_{c_{n}}$, and $\sigma^{\prime\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime\prime}_{c_{n}}$, where $\sigma^{\prime\prime}_{c_{i}}=\sigma_{c_{i}}$ or $\sigma^{\prime\prime}_{c_{i}}=\sigma^{\prime}_{c_{i}}$ for every $1\leq i\leq n$.
Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{m}}(s,\sigma^{\prime}_{c_{m}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})$, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})$. Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
(3) $\Rightarrow$ (2)
Assume (3). Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$.
First, we show (2a). Let $a\in\mathtt{AG}$. We want to show $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
  • Let $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$. Then, there is $\sigma_{a}\in\mathtt{JA}_{a}$ such that $t\in\mathsf{out}_{a}(s,\sigma_{a})$. By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$. By (3a), $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
  • Let $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. By (3a), $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$. Then, there is $\sigma_{\mathtt{AG}}$ such that $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. By (3a), $t\in\mathsf{out}_{a}(s,\sigma_{\mathtt{AG}}|_{a})$. Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$.
Second, we show (2b). We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. By (3a), $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Note $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ and $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ but $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
Note $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. We claim $\sigma_{\mathrm{C}}\nsubseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$. Assume $\sigma_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. Note $\mathrm{C}\cap\mathrm{D}=\emptyset$. Then $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. There is a contradiction. Similarly, we can show $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$.
Let $\sigma^{\prime}_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma^{\prime}_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(s)$ such that $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$. By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$ and $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}})$. Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ is a fusion of $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}$. By (3b), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. There is a contradiction.
Theorem 4.
Every single-coalition-first action model is a grand-coalition-first action model, but not vice versa.
The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counter-example.
Remarks
Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1]. Convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
5.2 Single-coalition-first neighborhood models
[…]
Definition 16 (Single-coalition-first neighborhood models).
A single-coalition-first neighborhood model is a tuple $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST})-\{\emptyset\})$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a general cover of $\mathtt{suc}(s)$.
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
Definition 17 (Neighborhood functions).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the neighborhood function $\mathtt{nei}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathrm{C}=\emptyset\text{ and }\mathtt{suc}(s)\neq\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathrm{C}\neq\emptyset\end{cases}$
Fact 6.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
For all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
Proof.
Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
Assume $\mathtt{suc}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is clearly a general cover of $\mathtt{suc}(s)$.
Assume $\mathrm{C}=\emptyset$ and $\mathtt{suc}(s)\neq\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is clearly a general cover of $\mathtt{suc}(s)$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Let $t\in\mathtt{suc}(s)$. Then there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in X_{a_{1}}$, …, there is $X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in X_{a_{n}}$. Then $t\in X_{a_{1}}\cap\dots\cap X_{a_{n}}$. Note $X_{a_{1}}\cap\dots\cap X_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
Definition 18 (Seriality, independence, and determinism of single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say:
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if and only if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if and only if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if and only if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
As above, we let the eight strings in the set $\mathrm{ES}$ correspond to the eight combinations of the three properties, respectively.
For any single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathtt{X}\in\mathrm{ES}$, we say $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model if it has the properties corresponding to $\mathtt{X}$.
Remarks
Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems. Models based on so-called tight alternating transition systems are single-coalition-first neighborhood $\mathtt{SID}$-models.
5.3 Representation of single-coalition-first action models by single-coalition-first neighborhood models
Theorem 5.
Every single-coalition-first action model is $z$-representable by a single-coalition-first neighborhood model.
This result is easy to show, and we skip its proof.
Theorem 6.
Every single-coalition-first neighborhood model $z$-represents a single-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood $\mathtt{X}$-model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as follows:
  • $\mathtt{AC}=\{\alpha_{a-s-X}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{ and }X\in\mathtt{nei}_{a}(s)\}$.
  • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-X}\in\mathtt{AC}$:
    $\mathsf{out}_{a}(s,\alpha_{x-u-X})=\begin{cases}X&\text{if $a=x$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
Let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
We claim that $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$.
Let $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. It suffices to show that $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathrm{C}=\emptyset$.
Then, $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Note:
$\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{n}})$
$\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}$
Let $X\in\mathtt{AE}(s,\mathrm{C})$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $X=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, where $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}},\dots,\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$. Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}(s,\sigma_{a_{n}})$ and $X$ is not empty. Then, $\sigma_{a_{1}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{a_{n}}\in\mathtt{av}_{a_{n}}(s)$.
Let $1\leq i\leq n$ and $\sigma_{a_{i}}=(a_{i},\alpha_{x-u-X_{i}})$. Note $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})\neq\emptyset$. Then, $\mathsf{out}(s,\sigma_{a_{i}})=X_{i}$, $\sigma_{a_{i}}=(a_{i},\alpha_{a_{i}-s-X_{i}})$ and $X_{i}\in\mathtt{nei}_{a_{i}}(s)$. Then, $X_{1}\cap\dots\cap X_{n}=X$. Then, $X\in\mathtt{nei}(s,\mathrm{C})$.
Let $X\in\mathtt{nei}(s,\mathrm{C})$. Then, there are $X_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{n}\in\mathtt{nei}_{a_{n}}(s)$ such that $X_{1}\cap\dots\cap X_{n}=X$. Then, $\alpha_{a_{1}-s-X_{1}}\in\mathtt{AC},\dots,\alpha_{a_{n}-s-X_{n}}\in\mathtt{AC}$. Let $\sigma_{a_{1}}=\{(a_{1},\alpha_{a_{1}-s-X_{1}})\},\dots,\sigma_{a_{n}}=\{(a_{n},\alpha_{a_{n}-s-X_{n}})\}$. Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})=X_{1},\dots,\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})=X_{n}$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=X_{1}\cap\dots\cap X_{n}=X$. Note $X$ is not empty. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$. Then, $X\in\mathtt{AE}_{\mathrm{C}}(s)$.
Theorem 7.
For every single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, if $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$, then for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
First, it is easy to see the following statements are equivalent:
  • $\mathtt{S}\textbf{-}\mathtt{AM}$ is serial;
  • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
  • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}(s)\neq\emptyset$;
  • for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$;
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial.
Second, it is easy to check the following statements are equivalent:
  • $\mathtt{S}\textbf{-}\mathtt{AM}$ is independent;
  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset,\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, and $\mathrm{Y}_{1},\mathrm{Y}_{2}\subseteq\mathtt{ST}$, if $\mathrm{Y}_{1}\in\mathtt{AE}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{AE}_{\mathrm{D}}(s)$, then $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$;
  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, and $\mathrm{Y}_{1},\mathrm{Y}_{2}\subseteq\mathtt{ST}$, if $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, then $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$;
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent.
Third, it is easy to see the following statements are equivalent:
  • $\mathtt{S}\textbf{-}\mathtt{AM}$ is deterministic;
  • for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})$ is a singleton;
  • for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{AE}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton;
  • for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton;
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic.
Then, it follows that for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
Theorem 8 (Representation of single-coalition-first action models by single-coalition-first neighborhood models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of single-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of single-coalition-first neighborhood $\mathtt{X}$-models.
Proof.
Let $\mathtt{X}\in\mathrm{ES}$.
Let $\mathtt{S}\textbf{-}\mathtt{AM}$ be a single-coalition-first action $\mathtt{X}$-model. By Theorem 5, $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{NM}$ is a single-coalition-first neighborhood $\mathtt{X}$-model.
Let $\mathtt{S}\textbf{-}\mathtt{NM}$ be a single-coalition-first neighborhood $\mathtt{X}$-model. By Theorem 6, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a single-coalition-first action $\mathtt{X}$-model.
6 Clear grand-coalition-first action models and clear single-coalition-first neighborhood models
6.1 Clear grand-coalition-first action models
Definition 19 (Clear grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Theorem 9.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, and let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
The following three conditions are equivalent:
  1. (1)
    for all $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, if $\sigma_{a}\neq\sigma^{\prime}_{a}$, then $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})=\emptyset$;
  2. (2)
    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$;
  3. (3)
    for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Assume (2) does not hold. Then, there is $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\neq\emptyset$.
Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
Then, there is $a\in\mathrm{C},\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\subseteq\sigma_{\mathrm{C}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathrm{C}}$ and $\sigma_{a}\neq\sigma^{\prime}_{a}$. Then, $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
Then, $w\in\mathsf{out}_{a}(s,\sigma_{a})$ and $w\in\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. There is a contradiction.
(2) $\Rightarrow$ (3)
(3) is a special case of (2).
(3) $\Rightarrow$ (1)
Assume (3). Assume (1) does not hold. Then, there is $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\neq\sigma^{\prime}_{a}$ and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$.
Let $w\in\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. There is a contradiction.
Theorem 10.
Every clear grand-coalition-first action model is a single-coalition-first action model, but not vice versa.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is essentially a single-coalition-first action model, we need to show:
  1. (1)
    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
  2. (2)
    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
  3. (3)
    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$,
    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
By Fact 5, (1) holds. Clearly, (3) holds. It remains to show (2).
Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
Assume $\mathrm{C}=\emptyset$. By the definition of $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathtt{suc}(s)$;
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
By Fact 4, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$. As $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, $\sigma^{1}_{\mathtt{AG}}=\dots=\sigma^{n}_{\mathtt{AG}}$. Then $\sigma_{\mathrm{C}}\subseteq\sigma^{1}_{\mathtt{AG}}$. Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
It is easy to find a single-coalition-first action model with a $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, $\sigma_{a}\neq\sigma^{\prime}_{a}$, and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. Then it is not a clear grand-coalition-first action model.
Remarks
Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models. As observed by Ågotnes and Alechina, the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
6.2 Clear single-coalition-first neighborhood models
Definition 20 (Clear single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$, and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
Theorem 11.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
The following two conditions are equivalent:
  1. (1)
    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
  2. (2)
    for all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
By Fact 6, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$. Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$. Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$. Without loss of any generality, let $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$. Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$.
Note $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$. Then $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Then, there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=X_{a_{1}}\cap\dots\cap X_{a_{n}}$, and there is $X^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=X^{\prime}_{a_{1}}\cap\dots\cap X^{\prime}_{a_{n}}$.
Then $w\notin X^{\prime}_{a_{i}}$ for some $1\leq i\leq n$. Note $w\in X_{a_{i}}$. Then $X_{a_{i}}\neq X^{\prime}_{a_{i}}$. Note $u\in X_{a_{i}}$ and $u\in X^{\prime}_{a_{i}}$. Then $X_{a_{i}}\cap X^{\prime}_{a_{i}}\neq\emptyset$. We have a contradiction.
(2) $\Rightarrow$ (1)
(1) is a special case of (2).
By 9, clear grand-coalition-first action models can be equivalently defined by three different ways, which respectively concern all agents, all coalitions and the grand coalition. It would be expected that the following condition is equivalent to the three conditions given in Theorem 11: for all $\mathtt{AG}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$. This is actually not the case. What follows is a counter-example.
Assume $\mathtt{AG}=\{a,b\}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, where:
  • $\mathtt{ST}=\{s_{0},s_{1},s_{2},s_{3}\}$;
  • $\mathtt{suc}(s_{0})=\{s_{1},s_{2},s_{3}\}$;
  • $\mathtt{nei}_{a}(s_{0})=\{\{s_{1},s_{2}\},\{s_{2},s_{3}\}\}$ and $\mathtt{nei}_{b}(s_{0})=\{\{s_{2}\},\{s_{1},s_{3}\}\}$.
Note $\mathtt{nei}_{a}(s_{0})$ is not a partition of $\mathtt{suc}(s_{0})$. However, it can be checked that $\mathtt{nei}_{\mathtt{AG}}(s_{0})=\{\{s_{1}\},\{s_{2}\},\{s_{3}\}\}$, which is a partition of $\mathtt{suc}(s_{0})$.
Remarks
Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems. It can be shown that those models are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
6.3 Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models
Theorem 12.
Every clear grand-coalition-first action model is $z$-representable by a clear single-coalition-first neighborhood model.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action model. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model. ∎
Theorem 13.
Every clear grand-coalition-first neighborhood model $z$-represents a clear single-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 8. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
From Theorem 12, Theorem 13, and Theorem 7, the following result follows:
Theorem 14 (Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of clear grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of clear single-coalition-first neighborhood $\mathtt{X}$-models.
7 Tree-like grand-coalition-first action models and tree-like single-coalition-first neigh-borhood models
7.1 Tree-like grand-coalition-first action models
Definition 21 (Histories in grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
Definition 22 (Tree-like grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history $\theta$ from $r$ to $s$.
Theorem 15.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
The following three conditions are equivalent:
  1. (1)
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$;
  2. (2)
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
  3. (3)
    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
Proof.
(1) $\Rightarrow$ (3)
Assume (1). Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $a\in\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is an $\mathtt{AG}$-history from $r$ to $r$. Assume $(r,\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{m},r)$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$. Then, $(r,\sigma_{\mathtt{AG}}^{1}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{m}|_{a},r)$ is an $a$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$. Let $(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$ be the $a$-history from $r$ to $s$. For every $1\leq i\leq n$, let $\sigma^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{a}\subseteq\sigma^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{i}_{\mathtt{AG}})$. Then $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $s$.
    Let $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ be an $\mathtt{AG}$-history from $r$ to $s$. We claim $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},$ $s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are identical.
    Assume $n\neq m$. Let $a\in\mathtt{AG}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then they are different. We have a contradiction.
    Assume $\sigma^{i}_{\mathtt{AG}}\neq\lambda^{i}_{\mathtt{AG}}$ for some $1\leq i\leq n$. Then there is $a\in\mathtt{AG}$ such that $\sigma^{i}_{\mathtt{AG}}|_{a}\neq\lambda^{i}_{\mathtt{AG}}|_{a}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then, they are different. We have a contradiction.
    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$. Let $a\in\mathtt{AG}$. Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$. Then, they are different. We have a contradiction.
(3) $\Rightarrow$ (2)
Assume (3). Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$. For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then $(s_{0},\lambda^{1}_{\mathtt{AG}},s_{1},\dots,\lambda_{\mathtt{AG}}^{m},s_{m})$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$. Let $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ be the $\mathtt{AG}$-history from $r$ to $s$. Then, $,(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},$$\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is an $\mathrm{C}$-history from $r$ to $s$.
    Assume $(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$ which is different from $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},$ $s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$.
    For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
    Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is different from $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. We have a contradiction.
(2) $\Rightarrow$ (1)
It is easy to see that (1) is a special case of (2).
Theorem 16.
Every tree-like grand-coalition-first action model is a clear grand-coalition-first action model, but not vice versa.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model. We are done if we can show for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Let $s\in\mathtt{ST}$, $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$. Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. Let $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, there are at least two paths $(s,\sigma_{\mathtt{AG}},w)$ and $(s,\sigma^{\prime}_{\mathtt{AG}},w)$ from $s$ to $w$. Then, there are at least two paths $(r,\dots,s,\sigma_{\mathtt{AG}},w)$ and $(r,\dots,s,\sigma^{\prime}_{\mathtt{AG}},w)$ from r to w. Note that $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like grand-coalition-first action model. Then, there is a contradiction.
It is easy to find a clear grand-coalition-first action model with a loop. Then, it is not tree-like.
Remarks
In the literature, Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models. They are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined here. It is required in their definition that for every state, there is a unique computation, which is a sequence of states, from the root to the state. Consequently, they are not clear.
7.2 Tree-like single-coalition-first neighborhood models
Definition 23 (Histories in single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
Definition 24 (Tree-like single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}$ $\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ from $r$ to $s$.
Theorem 17.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
The following two conditions are equivalent:
  1. (1)
    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$;
  2. (2)
    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
Proof.
(1) $\Rightarrow$ (2)
Assume (1). Let $r$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$. For every $1\leq i\leq m$, let $Z_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $Z_{i}\subseteq\mathrm{Y}_{i}$ and $s_{1}\in Z_{i}$. Then, $(s_{0},Z_{1},s_{1},\dots,Z_{m},s_{m})$ is an $a$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
  • Assume $s\neq r$.
    Assume $\mathrm{C}=\emptyset$. Let $a\in\mathtt{AG}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$. It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$ is a unique $a$-history from $r$ to $s$.
    Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$.
    For every $1\leq i\leq n$, let $Z_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq Z_{i}$. Then, $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
    Let $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ be a $\mathrm{C}$-history from $r$ to $s$. We claim $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ and $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ are identical.
    For every $1\leq i\leq m$, let $E_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $D_{i}\subseteq E_{i}$. Then, $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$ is a $a$-history from $r$ to $s$.
    Assume $n\neq m$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
    Assume $Z_{i}\neq D_{i}$ for some $1\leq i\leq n$. Then, $Z_{i}\cap D_{i}=\emptyset$. Then, $s_{i}\neq t_{i}$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,$ $\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
(2) $\Rightarrow$ (1)
It easy to see that (1) is a special case of (2).
By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which are respectively concerned with all agents, all coalitions, and the grand coalition. It would be expected that the following condition is equivalent to the two conditions in Theorem 17: there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathtt{AG}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$. In fact, this is not the case. The counterexample given after Theorem 11 also works here.
7.3 Representation of tree-like grand-coalition-first action models by tee-like single-coalition-first neighborhood models
Theorem 18.
Every tree-like grand-coalition-first action model is $z$-representable by a tree-like single-coalition-first neighborhood model.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model. By Theorem 4, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model. Then, it is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model. ∎
Theorem 19.
Every tree-like grand-coalition-first neighborhood model $z$-represents a tree-like single-coalition-first action model.
Proof.
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 8. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.
From Theorem 18, Theorem 19, and Theorem 7, the following result follows:
Theorem 20 (Representation of tree-like grand-coalition-first action models by tree-like single-coalition-first neighborhood models).
For every $\mathtt{X}\in\mathrm{ES}$, the class of tree-like grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models.
8 Each of those eight coalition logics is determined by these six kinds of models, too
For every finite nonempty sequence $\chi$, we use $\mathsf{last}(\chi)$ to denote the last element of $\chi$.
Theorem 21.
Let $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model.
There is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ such that for all $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$.
Proof.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},$ $\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ as follows:
  • $\mathtt{ST}^{\prime}=\mathtt{ST}_{0}^{\prime}\cup\mathtt{ST}_{1}^{\prime}\cup\mathtt{ST}_{2}^{\prime}\cup\dots$, where:
    • $\mathtt{ST}_{0}^{\prime}=\{w\}$;
    • $\mathtt{ST}_{k+1}^{\prime}=\{\chi-\sigma_{\mathtt{AG}}-u\mid\chi\in\mathtt{ST}_{k}^{\prime},\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}},u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
  • for every $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$:
    $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})=\{y\in\mathtt{ST}^{\prime}\mid y=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
  • for every $\chi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\chi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\mathsf{last}(\chi))\}$.
Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is what we want.
We show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is an $\mathtt{X}$-model.
[…]
  • Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent. Let $\chi\in\mathtt{ST}^{\prime}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ and $\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{D}}(\chi)$. Then, $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}^{\prime}_{\mathrm{D}}(\chi,\sigma_{\mathrm{D}})\neq\emptyset$.
    Then, there is $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$. Similarly, we can show $\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(\mathsf{last}(\chi))$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi))$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Let $u^{\prime}\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
    Then $u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma^{\prime}_{\mathtt{AG}})$ for some $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
    Then $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma^{\prime}_{\mathtt{AG}})$.
    Then, $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. Then, $\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\chi)$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is independent.
  • Assume $(\mathtt{G}\textbf{-}\mathtt{AM},w)$ is deterministic. Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic. Then, there is $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\chi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ is not a singleton. Let $s,s^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ where $s\neq s^{\prime}$. Let $s=\chi-\sigma_{\mathtt{AG}}-u$ and $s^{\prime}=\chi-\sigma_{\mathtt{AG}}-u^{\prime}$ where $u\neq u^{\prime}$. Then, $u,u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, there is a contradiction. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is deterministic.
It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is a tree-like model.
We show for all $\chi$, and $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},w\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w\Vdash\phi$.
We put an induction on $\phi$. We consider only the cases $\phi=p$ and $\phi=\langle[\mathrm{C}]\rangle\psi$, and skip other cases.
  • $\phi$ is a proposition letter $p$.
    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash p$. Then, $p\in\mathsf{label}^{\prime}(\chi)$. Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$.
    $\Rightarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash p$. Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$. Then, $p\in\mathsf{label}^{\prime}(\chi)$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$.
  • $\phi$ is $\langle[\mathrm{C}]\rangle\psi$.
    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ such that for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
    We are done if we can show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$. Let $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$.
    Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\text{and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$.
    Then, $y\in\mathsf{out}^{\prime}(\chi,\sigma_{\mathrm{C}})$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$. By the inductive hyphothesis, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(y)\Vdash\psi$. That is $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    $\Rightarrow$
    Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
    Note $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$. Then $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$. Then $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$ is not empty. Then $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$.
    Let $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})\text{ and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then, there is $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ such that $t=\mathsf{last}(y)$.
    Then, $t\in\mathsf{out}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
Theorem 22.
For every $\mathtt{X}\in\mathrm{ES}$ and $\phi\in\Phi$, the following statements are equivalent:
  1. (1)
    $\phi$ is valid with respect to the class of grand-coalition-first action $\mathtt{X}$-models;
  2. (2)
    $\phi$ is valid with respect to the class of single-coalition-first action $\mathtt{X}$-models;
  3. (3)
    $\phi$ is valid with respect to the class of single-coalition-first neighborhood $\mathtt{X}$-models;
  4. (4)
    $\phi$ is valid with respect to the class of clear grand-coalition-first action $\mathtt{X}$-models;
  5. (5)
    $\phi$ is valid with respect to the class of clear single-coalition-first neighborhood $\mathtt{X}$-models;
  6. (6)
    $\phi$ is valid with respect to the class of tree-like grand-coalition-first action $\mathtt{X}$-models;
  7. (7)
    $\phi$ is valid with respect to the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models.
Proof.
By Theorem 8 and Theorem 1, (2), and (3) are equivalent.
By Theorem 14 and Theorem 1, (4), and (5) are equivalent.
By Theorem 20 and Theorem 1, (6), and (7) are equivalent.
It suffices to show that (1), (2), (4) and (6) are equivalent. By Theorem 4, (1) implies (2). By Theorem 10, (2) implies (4). By Theorem 16, (4) implies (6). Suppose (6). Assume (1) does not hold. Then $\phi$ is false at a pointed grand-coalition-first action model $(\mathtt{G}\textbf{-}\mathtt{AM},w)$. Then, by Theorem 21, there is a tree-like grand-coalition-first model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ such that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)\nVdash\phi$. We have a contradiction. Then, (1).
9 Concluding remarks
Grand-coalition-first action models are those action models in which the outcome function for the grand coalition determines the outcome functions for other coalitions. Single-coalition-first action models are those action models in which the outcome functions for single coalitions determine the outcome functions for other coalitions. This class of models can be $z$-represented by the class of single-coalition-first neighborhood models. Clear grand-coalition-first action models are those single-coalition-first action models where different joint actions of a coalition have different outcomes. This class of models can be $z$-represented by the class of clear single-coalition-first neighborhood models. Tree-like single-coalition-first action models are those single-coalition-first action models where every state has a determined history. This class of models can be $z$-represented by the class of tree-like single-coalition-first neighborhood models. These kinds of models have different motivations.
Every kind of model has eight subclasses, determined by which of the properties seriality, independence of agents, and determinism hold. In this paper, we show that the subclasses of models that meet the same properties determine the same coalition logic.
Acknowledgments
Thanks go to Thomas Ågotnes, Valentin Goranko, and the audience of a logic seminar at Beijing Normal University.

### v2

@@
 In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism.
-They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions they meet.
-In this paper, we show that each of the eight coalition logics is also determined by the following six kinds of models, with the respective properties: single-coalition-first action models; single-coalition-first neighborhood models; clear grand-coalition-first action models; clear single-coalition-first neighborhood models; tree-like grand-coalition-first action models; tree-like single-coalition-first neighborhood models.
+They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions are met.
+In this paper, we show that each of the eight logics is also determined by six other kinds of models, with respective properties, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
 Keywords: coalition logics, action models, neighborhood models, seriality, independence of agents, determinism
@@
 1.1 Eight coalition logics determined by eight classes of general concurrent game models
-Coalition Logic $\mathsf{CL}$ proposed by Pauly ([10, 11]) is a central logic in logical research on strategic reasoning.
+Coalition Logic $\mathsf{CL}$, proposed by Pauly ([10, 11]), is a central logic in logical research on strategic reasoning.
 The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{C}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{C}$ ensures $\phi$.
@@
 The first kind is concurrent game models.
-Roughly, in a concurrent game model: there are some states; there are some agents who form coalitions; for every coalition, there is an availability function, which specifies available joint actions of the coalition at states; for every coalition, there is an outcome function, which specifies possible outcome states of joint actions of the coalition.
-The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action in the model.
+Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; for every coalition, there is an availability function, specifying available joint actions of the coalition at states; for every coalition, there is an outcome function, specifying possible outcome states of joint actions of the coalition.
+The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
 The second kind is alpha neighborhood models.
-Roughly, in an alpha neighborhood model: there are some states; there are some agents who form coalitions; for every coalition, there is an alpha neighborhood function, which specifies the alpha powers of the coalition.
-An alpha power is a scope into which the coalition can force the next moment to fall.
+Alpha neighborhood models do not have actions.
+Consequently, they do not have availability functions and outcome functions.
+Instead, for every coalition, there is a neighborhood function that specifies the alpha powers of the coalition.
+An alpha power is a set of states into which the coalition can force the next moment to fall.
 The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
-[9] proposed that concurrent game models have three too-strong assumptions: the independence of agents, seriality, and determinism.
-Based on general concurrent game models, which do not have the three assumptions, a Minimal Coalition Logic $\mathsf{MCL}$ was presented.
-Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some special kinds of situations.
+[9] argued that concurrent game models have three too strong assumptions: seriality, the independence of agents, and determinism.
+Based on general concurrent game models, which do not have the three assumptions, they proposed a Minimal Coalition Logic $\mathsf{MCL}$.
+Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some kinds of situations.
 Considering which of the three properties we want to keep, there are eight coalition logics in total.
 [8] showed the completeness of the eight logics in a uniform way.
-It should be pointed out that the completeness $\mathsf{CL}$ is already given in [11, 5] and the completeness of $\mathsf{MCL}$ is already given in [9].
-General concurrent game models are called grand-coalition-first action models in this paper.
 1.2 Our work
-In this work, we show that each of the eight coalition logics is also determined by six other kinds of models with respective properties.
+In this paper, we call general concurrent game models grand-coalition-first action models.
+In this work, we show that each of the eight coalition logics is also determined by six other kinds of models, with respective properties.
 The first kind is single-coalition-first action models.
-In this kind of models, the outcome function for single coalitions determines the outcome functions for other coalitions and the availability functions for all coalitions.
-    This kind of models makes good sense for situations where we know the space of future possibilities and how agents’ behavior changes the future beforehand.
+In this kind of models, the outcome functions for single coalitions determine the outcome functions for other coalitions.
+This kind of models makes good sense for situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other.
 The second kind is single-coalition-first neighborhood models.
 In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition.
-An actual power is a minimal scope in which the coalition can force the next moment to fall.
+An actual power is a minimal scope into which the coalition can force the next moment to fall.
 In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
-  • The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition do not have the same outcome.
-  • The fourth kind is clear single-coalition-first neighborhood models. In this kind of models, the powers of an agent are pairwise disjoint.
-  • The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history.
-  • The sixth kind is tree-like single-coalition-first neighborhood models. In this kind of models, every state has a determined history, too.
+The third kind is clear grand-coalition-first action models.
+In this kind of models, different joint actions of the grand coalition have different outcomes.
+Clear grand-coalition-first action models are quite common in game theory.
+The fourth kind is clear single-coalition-first neighborhood models.
+In this kind of models, the actual powers of an agent are pairwise disjoint.
+The fifth kind is tree-like grand-coalition-first action models.
+In this kind of models, every state has a determined history.
+The sixth kind is tree-like single-coalition-first neighborhood models.
+In this kind of models, every state has a determined history, too.
 The rest of the paper is structured as follows.
-  • In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from the former to the latter.
+In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from action semantics to neighborhood semantics.
 In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models.
-  • In Section 4, we discuss compositionality of availability functions and outcome functions in action models.
-  • In Section 5: we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models, and show they can represent single-coalition-first action models.
-  • In Section 6: we define clear grand-coalition-first action models, and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models, and show that they can represent clear grand-coalition-first action models.
-  • In Section 7, we define tree-like grand-coalition-first action models, and show they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models, and show they can represent tree-like grand-coalition-first action models.
+In Section 4, we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models and show that they can represent single-coalition-first action models.
+In Section 5, we define clear grand-coalition-first action models and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models and show that they can represent clear grand-coalition-first action models.
+In Section 6, we define tree-like grand-coalition-first action models and show that they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models and show that they can represent tree-like grand-coalition-first action models.
 In Section 7, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties.
-  • We make some concluding remarks in Section 9.
+We conclude in Section 8.
 2 General settings of coalition logics
 2.1 Language
-Let $\mathtt{AP}$ be a countable set of atomic propositions and $\mathtt{AG}$ be a finite nonempty set of agents.
-The subsets of $\mathtt{AG}$ are called coalitions.
-$\mathtt{AG}$ is called the grand coalition.
+Let $\mathtt{AP}$ be a countable set of atomic propositions, and $\mathtt{AG}$ be a finite nonempty set of agents.
+The subsets of $\mathtt{AG}$ are called coalitions, and $\mathtt{AG}$ is called the grand coalition.
+In the sequel, when no confusion arises, we often write $a$ instead of $\{a\}$, where $a\in\mathtt{AG}$.
 Definition 1 (The language $\Phi$).
@@
 $\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
-The formula $\langle[\mathrm{C}]\rangle\phi$ indicates that some available joint action of $\mathrm{C}$ ensures $\phi$.
+The formula $\langle[\mathrm{C}]\rangle\phi$ indicates some available joint action of $\mathrm{C}$ ensures $\phi$.
 The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual.
-Define the dual $[\langle\mathrm{C}\rangle]\phi$ of $\langle[\mathrm{C}]\rangle\phi$ as $\neg\langle[\mathrm{C}]\rangle\neg\phi$, indicating every available joint action of $\mathrm{C}$ enables $\phi$.
 2.2 Action semantics
 Let $\mathtt{AC}$ be a nonempty set of actions.
-For any $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is a set of joint actions of $\mathrm{C}$.
-For any $a\in\mathtt{AG}$, we write $\mathtt{JA}_{a}$ instead of $\mathtt{JA}_{\{a\}}$.
+For every $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is a set of joint actions of $\mathrm{C}$.
+Joint actions of the grand coalition are called action profiles.
+In the sequel, we sometimes use sequences of actions to represent joint actions of coalitions, where an order among agents is implicitly presupposed.
 We define $\mathtt{JA}=\bigcup\{\mathtt{JA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
+For every $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{D}\subseteq\mathrm{C}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we use $\sigma_{\mathrm{C}}|_{\mathrm{D}}$ to indicate the restriction of $\sigma_{\mathrm{C}}$ to $\mathrm{D}$, which is in $\mathtt{JA}_{\mathrm{D}}$.
 Definition 2 (Action models).
@@
   • $\mathtt{AC}$ is a nonempty set of actions.
-  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an available action function for $\mathrm{C}$.
-  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
+  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an availability function for $\mathrm{C}$.
+  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
   • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Different logics might put different constraints on action models.
-Definition 3 (Action semantics of $\Phi$).
-$\mathtt{AM},s\Vdash\top$ always
+Different coalition logics might put different constraints on action models.
+Definition 3 (Action semantics).
+$\mathtt{AM},s\Vdash\top$
 $\mathtt{AM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
@@
 $\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
-It can be verified:
-$\mathtt{AM},s\Vdash[\langle\mathrm{C}\rangle]\phi$Leftrightarrow for all $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s,\mathrm{C})$, there is $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ such that $\mathtt{AM},t\Vdash\phi$
 2.3 Neighborhood semantics
@@
   • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-A neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is called an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is closed under supersets.
-Different logics may put different constraints on neighborhood models.
-Definition 5 (Neighborhood semantics of $\Phi$).
-$\mathtt{NM},s\Vdash\top$ always
+A neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is closed under supersets.
+Different coalition logics might put different constraints on neighborhood models.
+Definition 5 (Neighborhood semantics).
+$\mathtt{NM},s\Vdash\top$
 $\mathtt{NM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
@@
 $\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
-The following fact indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing which formulas it satisfies.
+The following fact, which is easy to verify, indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing the set of formulas it satisfies.
 Fact 1.
-Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\mathtt{AG},$$\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
+Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
 Then, for every $s\in\mathtt{ST}$ and $\phi\in\Phi$, $\mathtt{NM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
 2.4 Transformation of action semantics to neighborhood semantics
-Definition 6 (Actual and alpha effectivity functions of action models).
-Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model, and $\mathrm{C}\subseteq\mathtt{AG}$.
-The function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
+Definition 6 (Actual effectivity functions and alpha effectivity functions of action models).
+Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
+For every $\mathrm{C}\subseteq\mathtt{AG}$, the function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ in $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
 $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
-The function $\mathtt{LE}_{\mathrm{C}}$ defined as follows is called the alpha effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
+For every $\mathrm{C}\subseteq\mathtt{AG}$, the function $\mathtt{LE}_{\mathrm{C}}$ defined as follows is called the alpha effectivity function for $\mathrm{C}$ in $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
 $\mathtt{LE}_{\mathrm{C}}(s)=\{\mathrm{Y}\subseteq\mathtt{ST}\mid\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathrm{Y}\text{ for some }\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
-In a sense, actual effectivity functions and alpha effectivity functions of action models contain important information on coalition powers.
 Actual effectivity functions are more transparent in the following sense: each element of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions.
 Definition 7 (Action models $z$-representable by neighborhood models and $\alpha$-representable by alpha neighborhood models).
-We say a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $z$-represents an action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$.
-We say a class of neighborhood models $\mathbf{NM}$ $z$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
-We say an alpha neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $\alpha$-represents an action model $\mathtt{AM}$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{LE}_{\mathrm{C}}$.
-We say a class of alpha neighborhood models $\mathbf{NM}$ $\alpha$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}$, and (2) every alpha neighborhood model in $\mathbf{NM}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
+An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $z$-representable by a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
+A class of action models $\mathbf{AM}$ is $z$-representable by a class of neighborhood models $\mathbf{NM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
+An action model $\mathtt{AM}$ is $\alpha$-representable by an alpha neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{LE}_{\mathrm{C}}=\mathtt{nei}^{\alpha}_{\mathrm{C}}$.
+A class of action models $\mathbf{AM}$ is $\alpha$-representable by a class of alpha neighborhood models $\mathbf{NM}^{\alpha}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}^{\alpha}$, and (2) every alpha neighborhood model in $\mathbf{NM}^{\alpha}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
 Intuitively, that a neighborhood model $\mathtt{NM}$ represents an action model $\mathtt{AM}$ means that $\mathtt{NM}$ contains all the information of $\mathtt{AM}$ to evaluate formulas of $\Phi$.
+The following result can be easily shown:
 Theorem 1 (Representability implies equivalence).
   1. (1)
-    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
-    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
+    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
+    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
   2. (2)
-    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
-    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
-Proof.
-Here, we show only (1). We can show (2) in a similar way.
-Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model such that $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$.
-Let $s\in\mathtt{ST}$ and $\phi\in\Phi$.
-We want to show $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
-The proof is by induction of $\phi$.
-It is trivial when $\phi$ is a proposition letter p, $\phi$ is $\top$, $\phi$ is $\neg\psi$ and $\phi$ is $\psi\wedge\theta$.
-We are done if we can show when $\phi$ is $\langle[\mathrm{C}]\rangle\psi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
-$\Rightarrow$
-Assume $\mathtt{AM},s\Vdash\phi$.
-Then, $\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\psi$.
-Then, there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$.
-Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$.
-Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
-Then, for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$.
-Then, $\mathtt{NM},s\Vdash\phi$.
-$\Leftarrow$
-Assume $\mathtt{NM},s\Vdash\phi$.
-Then, $\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\psi$.
-Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$.
-Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$.
-Then, $\mathrm{Y}\in\mathtt{AE}_{\mathrm{C}}(s)$.
-Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$.
-Then, $\mathtt{AM},s\Vdash\phi$.
+    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}^{\alpha}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}^{\alpha}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
+    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}^{\alpha}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}^{\alpha}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}^{\alpha}$.
 3 Eight coalition logics determined by eight classes of grand-coalition-first action models
@@
   • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Definition 9 (Outcome functions and availability functions).
+Definition 9 (Outcome functions and availability functions of grand-coalition-first action models).
 Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
@@
 $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
-Outcome functions in grand-coalition-first action models are given in a from-whole-to-part way: the outcome states of a coalition $\mathrm{C}$ performing a joint action $\sigma_{\mathrm{C}}$ at a state $s$ are determined by the outcome states of the grand coalition performing action profiles containing $\sigma_{\mathrm{C}}$ at $s$.
+Definition 10 (Successor functions of grand-coalition-first action models).
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},$ $\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
+Define the successor function $\mathtt{suc}$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ as follows: for all $s\in\mathtt{ST}$,
+$\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}.$
+For every set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$.
+A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$.
+It is easy to verify the following fact:
 Fact 2.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
-Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\mid\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ for some }\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(s)\}$.
-This fact indicates that availability functions in grand-coalition-first action models also are given in a from-whole-to-part way: available joint actions of a coalition $\mathrm{C}$ at a state are determined by available joint actions the grand coalition at the state.
-Definition 10 (Successor functions).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
-Define the successor function $\mathtt{suc}$ as follows: for all $s\in\mathtt{ST}$,
-$\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}.$
-For any set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$.
-Fact 3.
 Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
 Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathtt{suc}(s)$.
-Proof.
-$\Rightarrow$
-Let $\mathrm{C}\subseteq\mathtt{AG}$.
-Let $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$.
-Then, there is $\sigma_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
-Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$.
-Then, $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$.
-That is $w\in\mathtt{suc}(s)$.
-$\Leftarrow$
-Let $w\in\mathtt{suc}(s)$.
-Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$.
-Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$.
-Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
-Then, $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$.
+Fact 3.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+Then:
+  1. (1)
+    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
+  2. (2)
+    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})$.
+Note that neither of the converses of the two statements holds. Figure 1 indicates a counterexample for them.
+$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
+Figure 1: This figure is used to show that in grand-coalition-first action models, the set of outcome states of a joint action might not contain the intersection of the sets of outcome states of the joint action’s parts.
+A grand-coalition-first action model is indicated in the figure, where we suppose $\mathtt{AG}=\{a,b\}$.
+It can be seen: (1) $\mathsf{out}_{a}(s_{0},\alpha_{1})$, the set of outcome states of $a$ performing $\alpha_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, (2) $\mathsf{out}_{a}(s_{0},\beta_{1})$, the set of outcome states of $b$ performing $\beta_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, but (3) $\mathsf{out}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$, the set of outcome states of $a$ and $b$ respectively performing $\alpha_{1}$ and $\beta_{1}$ at $s_{0}$, does not contain the intersection of $\mathsf{out}_{a}(s_{0},\alpha_{1})$ and $\mathsf{out}_{b}(s_{0},\beta_{1})$.
 Definition 11 (Seriality, independence, and determinism of grand-coalition-first action models).
@@
   • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
-  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
+  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
   • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, ${\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})}$ is a singleton.
-We use $\mathrm{ES}$ to indicate the set of the eight strings: $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$.
-We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively.
-We use the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ to correspond to the eight combinations of the three properties, respectively.
-For every $\mathtt{X}\in\mathrm{ES}$ and grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$, we say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action $\mathtt{X}$-model if $\mathtt{G}\textbf{-}\mathtt{AM}$ has the properties corresponding to $\mathtt{X}$.
+We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively, and let the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ correspond to the eight combinations of the three properties, respectively.
+We use $\mathrm{ES}$ to indicate the set of the eight strings.
+For every $\mathtt{X}\in\mathrm{ES}$ and grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$, we say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a $\mathtt{X}$-model if $\mathtt{G}\textbf{-}\mathtt{AM}$ has the properties corresponding to $\mathtt{X}$.
+For every $\mathtt{X}\in\mathrm{ES}$, we use $\mathsf{MCL+X}$ to refer to the coalition logic determined by the class of grand-coalition-first action $\mathtt{X}$-models.
 3.2 Eight axiomatic systems
@@
 Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
-Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$
-where $\mathrm{C}\cap\mathrm{D}=\emptyset$
+Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
 Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
 Definition 13 (Axiomatic systems for $\mathsf{MCL+X}$).
-For all $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inferences of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
+For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inference rules of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
 Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [8]).
-The axiomatic system for $\mathsf{MCL+X}$ given in Definition 13 is sound and complete with respect to the set of valid formulas in $\Phi$ relative to the class of grand-coalition-first action $\mathtt{X}$-models.
+For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ given in Definition 13 is sound and complete with respect to the set of valid formulas of $\mathsf{MCL+X}$.
 Remarks
@@
 Are the eight classes of grand-coalition-first action models $z$-representable?
-It is not known yet.
-4 About compositionality of availability functions and outcome functions in grand-coalition-first action models
-We say availability functions of a grand-coalition-first action model is compositional if the availability function of a coalition can be derived from the availability functions of its members in a natural way.
-In a similar sense, we talk about compositionality of outcome functions of grand-coalition-first action models.
-4.1 Neither outcome functions nor availability functions in grand-coa-lition-first action models are compositional
-The following two facts are easy to verify, and we skip their proofs.
-Fact 4.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
-Then:
-  1. (1)
-    1. (a)
-      for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
-    2. (b)
-      for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})$.
-  2. (2)
-    1. (a)
-      for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
-    2. (b)
-      for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, then $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
-Note that none of the converses of these items holds. Figure 1 indicates a counter-example for the first two items and Figure 2 indicates a counter-example for the last two items.
-$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{1}\atop\alpha_{2}\beta_{2}$$\alpha_{1}\beta_{2}\atop\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
-Figure 1: An example of that outcome functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
-$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{2}$$\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
-Figure 2: An example of that availability functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
-4.2 Availability functions in grand-coalition-first action models meeting independence of agents are compositional
-The following fact is easy to check:
-Fact 5.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action $\mathtt{I}$-model.
-Then:
-  1. (1)
-    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$: $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$ if and only if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
-  2. (2)
-    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$: $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ if and only if $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
-Remarks
-Models of Coalition Logic are grand-coalition-first action $\mathtt{SID}$-models.
-In those models, availability functions are compositional, but outcome functions might not be.
-Figure 1 indicates a counter-example.
+All this is yet unknown.
+4 Single-coalition-first action models and single-coalition-first neighborhood models
 4.1 Outcome functions are compositional in many situations
+We say outcome functions of a class of action models are compositional if there is a way such that for every action model in the class, the outcome function of a coalition is determined by the outcome functions of its members in this way.
+Outcome functions in grand-coalition-first action models are not compositional. Figure 2 gives a counterexample for this.
+$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{1}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
+$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{2}$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
+Figure 2: This figure offers a counterexample that outcome functions in grand-coalition-first action models are not generally compositional.
+We suppose $\mathtt{AG}=\{a,b\}$.
+Two grand-coalition-first action models $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$ are depicted above, where $\{\mathsf{out}^{1}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\{\mathsf{out}^{2}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$.
+In $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{1}\}$ In $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$: $\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{2}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{2}\}$.
+Note: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\mathsf{out}^{2}_{b}(s_{0},\beta_{1})$, but $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))\neq\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$.
+This implies that outcome functions in grand-coalition-first action models are not generally compositional: there is no way such that for every grand-coalition-first action model, the outcome function of a coalition is determined by the outcome functions of its members in this way.
 There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other.
-Outcome functions are compositional in these situations.
+In these situations, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
 What follows are two examples, and the second is from [3].
 Example 1.
-Adam and Bob control a ship lock, which has two doors.
-Adam can close or open the front door and Bob can close or open the back door.
-They cannot open the two doors at the same time.
-This scenario can be represented by the grand-coalition-first action model depicted in Figure 3.
-$s_{1}$$s_{2}$$s_{3}$$\{c_{f},c_{b}\}$$\{c_{f}\}$$\{c_{b}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{open_{b}})$$(\mathtt{skip},\mathtt{close_{b}})$$(\mathtt{open_{f}},\mathtt{skip})$$(\mathtt{close_{f}},\mathtt{skip})$
-Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1, where outcome functions are compositional.
-Here $c_{f}$ and $c_{b}$, respectively, express the propositions the front door is closed and the back door is closed.
-Here are some instances of outcome functions: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{2},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{2},\mathtt{skip})=\{s_{2}\}\cap\{s_{1},s_{2}\}=\{s_{2}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
+Adam and Bob control a ship lock with two doors: a front door and a back door.
+Adam can make the front door closed or open, Bob can make the back door closed or open, but they cannot make the two doors open at the same time.
+This situation can be represented by the grand-coalition-first action model depicted in Figure 3, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
+$s_{1}$$s_{2}$$s_{3}$$\{c_{f},c_{b}\}$$\{c_{f}\}$$\{c_{b}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{open}\text{-}\mathtt{b})$$(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b})$$(\mathtt{open}\text{-}\mathtt{f},\mathtt{skip})$$(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip})$
+Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1.
+Here, $c_{f}$ and $c_{b}$, respectively, express the front door is closed and the back door is closed.
+It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
+For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{close}\text{-}\mathtt{f})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close}\text{-}\mathtt{f})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
 Example 2.
 A system has two processes, $a$ and $b$.
-The process $a$ can assign values $1$ or 00 to the variable $x$, and the process $a$ can assign values $1$ or 00 to the variable $y$.
-When $x=0$, then $a$ can leave the value of $x$ unchanged or change it to $1$.
-When $x=1$, then $a$ leaves the value of $x$ unchanged.
-Similarly, when $y=0$, then $b$ can leave the value of $y$ unchanged or change it to $1$.
-When $y=1$, then $b$ leaves the value of $y$ unchanged.
-Figure 4 indicates a grand-coalition-first model for this scenario, where outcome functions are compositional.
+The process $a$ can assign $1$ or 00 to the variable $x$, and the process $b$ can assign $1$ or 00 to the variable $y$.
+When $x=0$, $a$ can leave the value of $x$ unchanged or change it to $1$.
+When $x=1$, $a$ leaves the value of $x$ unchanged.
+Similarly, when $y=0$, $b$ can leave the value of $y$ unchanged or change it to $1$.
+When $y=1$, $b$ leaves the value of $y$ unchanged.
+Figure 4 indicates a grand-coalition-first model for this scenario, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
 $s_{1}$$s_{2}$$s_{3}$$s_{4}$$\{x_{1},y_{1}\}$$\{x_{1},y_{0}\}$$\{x_{0},y_{1}\}$$\{x_{0},y_{0}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{x:=1},\mathtt{y:=1})$$(\mathtt{x:=1},\mathtt{skip})$$(\mathtt{skip},\mathtt{y:=1})$
-Figure 4: This figure indicates a grand-coalition-first action model where outcome functions are compositional.
-Here are some instances: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
-The following settings about agency can be found in the literature, such as [6] and [7]: there are some agents; every atomic proposition is controlled by at most one agent.
+Figure 4: This figure indicates a grand-coalition-first action model for the situation of Example 2.
+It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
+For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\mathtt{AG}}(s_{4},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{4},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{4},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
+We want to mention that some literature, such as [6] and [7], deals with agency by propositional control: there are some agents; every atomic proposition is controlled by at most one agent.
 In these settings, outcome functions are compositional.
-5 Single-coalition-first action models and single-coalition-first neighborhood models
 4.2 Single-coalition-first action models
@@
   • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Definition 15 (Outcome functions and availability functions).
+Definition 15 (Outcome functions and availability functions of single-coalition-first action models).
 Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model.
@@
 For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
-    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
+$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}$
 Clearly, the outcome functions in single-coalition-first action models are compositional.
-Let $\mathtt{AC}$ be a set of actions and $\mathtt{JA}_{\mathrm{C}}$ be the set of joint actions of $\mathrm{C}$.
-For any $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}$, if for any $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma^{\prime}_{\mathrm{C}}|_{a}$.
+For all $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, if for all $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}$ equals to $\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime}_{\mathrm{C}}|_{a}$.
+The following result offers two alternative definitions of sing-coalition-first action models.
 Theorem 3.
-Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
-The following three conditions are equivalent:
+Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $s\in\mathtt{ST}$.
+The following three sets of conditions are equivalent:
   1. (1)
-    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
     1. (a)
@@
     2. (b)
-      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$, where $\mathrm{C}\neq\emptyset$.
+      for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
+    It is easy to check this set of conditions is equivalent to the set of constraints on single-coalition-first action models given in Definition 14.
   2. (2)
-    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$:
     1. (a)
@@
     2. (b)
-      $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+      for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
   3. (3)
-    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
     1. (a)
-      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$;
+      for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$;
     2. (b)
-      for all fusion $\sigma^{\prime\prime}_{\mathrm{C}}$ of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, for all $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$, then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
+      for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma_{\mathrm{C}}^{\prime},\sigma_{\mathrm{C}}^{\prime\prime}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
 Proof.
@@
 Assume (1).
-Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$.
-It suffices to show (2b).
-Assume $\mathrm{C}=\emptyset$.
-Then, $\sigma_{\mathrm{C}}=\emptyset$.
-Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{\mathrm{D}}$.
-Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+We only need to show (2b).
+Let $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$.
+We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}=\emptyset$.
+Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\emptyset}(s,\emptyset)=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}\neq\emptyset$.
+Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+It suffices to show $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\emptyset}(s,\emptyset)$.
 Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
-By (1b), $t\in\mathsf{out}_{b}(s,\sigma_{b})$, where $b\in\mathrm{D}$.
-Then, $t\in\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}$.
-Note $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
+Let $a\in\mathrm{D}$.
+By (1b), $t\in\mathsf{out}_{a}(s,\sigma_{\mathrm{D}}|_{a})$.
+Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
 Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
@@
 Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
-Assume $\mathrm{D}=\emptyset$.
-Similarly, we can show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+The case that $\mathrm{C}\neq\emptyset$ and $\mathrm{D}=\emptyset$ can be handled similarly.
 Assume $\mathrm{C}\neq\emptyset$ and $\mathrm{D}\neq\emptyset$.
-Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\mathrm{D}=\{b_{1},\dots,b_{m}\}$.
-Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$ and $\sigma_{\mathrm{D}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$.
-By (1b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$ and $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$.
-Then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$.
-Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}\cup\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$.
-By (1b), $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$.
-Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, $\mathrm{D}=\{b_{1},\dots,b_{m}\}$, $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, and $\sigma_{\mathrm{D}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$.
+By (1b), $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
 (2) $\Rightarrow$ (1)
-Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. It suffices to show (1b).
-Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$.
-Note $\mathrm{C}=\{c_{1}\}\cup\mathrm{C}-\{c_{1}\}$.
-By (2b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ = $\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$.
-Note $\mathrm{C}-\{c_{1}\}=\{c_{2}\}\cup\mathrm{C}-\{c_{1},c_{2}\}$.
-By (2b), $\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$ = $\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2}\}})$.
-Finally, we get
-$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}}(s,$ $\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}})$.
-That is, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap$ $\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$.
-Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
+Assume (2).
+It suffices to show (1b).
+Let $\mathrm{C}$ be a nonempty subset of $\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
+Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
+Note $\mathrm{C}=\{a_{1}\}\cup\mathrm{C}-\{a_{1}\}$.
+By (2b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\mathsf{out}_{\mathrm{C}-\{a_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1}\}})$.
+Note $\mathrm{C}-\{a_{1}\}=\{a_{2}\}\cup\mathrm{C}-\{a_{1},a_{2}\}$.
+By (2b), $\mathsf{out}_{\mathrm{C}-\{a_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1}\}})=\mathsf{out}_{a_{2}}(s,\sigma_{a_{2}})\cap\mathsf{out}_{\mathrm{C}-\{a_{1},a_{2}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1},a_{2}\}})$.
+Finally, we have $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
 (2) $\Rightarrow$ (3)
-Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
+Assume (2).
 First, we show (3a).
+Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
+We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
 Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
 We want to show $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
-We are done if we can show there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$.
-It suffices to show that there is a joint action $\sigma_{\overline{\mathrm{C}}}$ of $\overline{\mathrm{C}}$ such that $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
-    Assume $\mathrm{C}=\emptyset$.
-Then, $\sigma_{\mathrm{C}}=\emptyset$.
-Then, $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
+It suffices to show there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$.
+By (2b), it suffices to show that there is $\sigma_{\overline{\mathrm{C}}}\in\mathtt{JA}_{\overline{\mathrm{C}}}$ such that $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
+By (1a) and (1b), it is easy to verify $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
+Assume ${\overline{\mathrm{C}}}=\emptyset$.
+Clearly, what we want holds.
+Assume ${\overline{\mathrm{C}}}\neq\emptyset$.
 Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$.
 Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
-Then, there are $\sigma_{a_{1}},\dots,\sigma_{a_{n}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}}),\dots,w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
+Then, there is $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})$, …, there is $\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$ such that $w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
 Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
-By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
-Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
-    Assume $\mathrm{C}\neq\emptyset$.
-Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
-Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
-Then, $w\in\bigcup\{\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\mid\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}\}$.
-Then, $w\in\mathtt{suc}(s)$.
-Assume ${\overline{\mathrm{C}}}=\emptyset$.
-Then, $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathtt{suc}(s)$.
-Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
-Assume ${\overline{\mathrm{C}}}\neq\emptyset$.
-Assume ${\overline{\mathrm{C}}}=\{b_{1},\dots,b_{n}\}$.
-As for every $b\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathtt{suc}(s)$, there are $\sigma_{b_{1}},\dots,\sigma_{b_{n}}$ such that $w\in\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}}),\dots,w\in\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$.
-Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{n}}$.
-By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$.
-Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
+By (1b), $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
 Let $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
-We want to show $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
 Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$.
 Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$.
-By (2b), $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
-Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
+By (2b), $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
 Second, we show (3b).
-Let $\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, and $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
-Assume $\mathrm{C}=\emptyset$.
-Then, $\sigma_{\mathrm{C}}=\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{\mathrm{C}}=\emptyset$.
-Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
-Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
+Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma_{\mathrm{C}}^{\prime},\sigma_{\mathrm{C}}^{\prime\prime}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$.
+We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
+Let $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
+Assume $\mathrm{C}=\emptyset$. It is easy to see $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
 Assume $\mathrm{C}\neq\emptyset$.
 Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$.
-Let $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$, where $\sigma_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma_{c_{n}}\in\mathtt{JA}_{c_{n}}$, $\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime}_{c_{n}}$, where $\sigma^{\prime}_{c_{1}}\in\mathtt{JA}_{c_{1}},\dots,\sigma^{\prime}_{c_{n}}\in\mathtt{JA}_{c_{n}}$, and $\sigma^{\prime\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime\prime}_{c_{n}}$, where $\sigma^{\prime\prime}_{c_{i}}=\sigma_{c_{i}}$ or $\sigma^{\prime\prime}_{c_{i}}=\sigma^{\prime}_{c_{i}}$ for every $1\leq i\leq n$.
-Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{m}}(s,\sigma^{\prime}_{c_{m}})$.
-Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})$, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime}_{c_{n}})$.
+Let $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$, $\sigma^{\prime}_{\mathrm{C}}=\sigma^{\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime}_{c_{n}}$, and $\sigma^{\prime\prime}_{\mathrm{C}}=\sigma^{\prime\prime}_{c_{1}}\cup\dots\cup\sigma^{\prime\prime}_{c_{n}}$, where $\sigma^{\prime\prime}_{c_{i}}=\sigma_{c_{i}}$ or $\sigma^{\prime\prime}_{c_{i}}=\sigma^{\prime}_{c_{i}}$ for every $i$ such that $1\leq i\leq n$.
+By (1b), $t\in\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})$, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime}_{c_{n}})$.
 Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})$, …, $t\in\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})$.
-Then, $t\in\mathsf{out}_{c_{1}}(s,\sigma^{\prime\prime}_{c_{1}})\cap\dots\cap\mathsf{out}_{c_{n}}(s,\sigma^{\prime\prime}_{c_{n}})=\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
+By (1b), $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
 (3) $\Rightarrow$ (2)
 Assume (3).
-Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$.
 First, we show (2a). Let $a\in\mathtt{AG}$. We want to show $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
@@
 By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$.
-Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$.
-By (3a), $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
+Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$, which equals to $\mathsf{out}_{\emptyset}(s,\emptyset)$ by (3a).
 Let $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
@@
 Second, we show (2b).
+Let $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$.
 We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
 Then, $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
-Assume $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ but $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
+Assume $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
 Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
-Note $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
+Assume $t\notin\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
+We want to get a contradiction.
 We claim $\sigma_{\mathrm{C}}\nsubseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$.
@@
 Then $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
-By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
-There is a contradiction.
+By (3a), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$, which is impossible.
 Similarly, we can show $\sigma_{\mathrm{D}}\nsubseteq\sigma_{\mathtt{AG}}$.
-Let $\sigma^{\prime}_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma^{\prime}_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(s)$ such that $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$.
+Let $\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$ such that $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$.
 Then, $\sigma^{\prime}_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$ and $\sigma_{\mathrm{C}}\cup\sigma^{\prime}_{\mathrm{D}}\subseteq\sigma_{\mathtt{AG}}$.
@@
 By (3b), $t\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
-There is a contradiction.
+We have a contradiction.
 Theorem 4.
 Every single-coalition-first action model is a grand-coalition-first action model, but not vice versa.
-The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counter-example.
+The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counterexample.
 Remarks
 Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1].
-Convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
+It can be verified convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
 4.3 Single-coalition-first neighborhood models
-[…]
+Let $\mathtt{ST}$ be a nonempty set of states and $\Delta_{1},\Delta_{2}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}$.
+Define $\Delta_{1}\odot\Delta_{2}=\{\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\mid\mathrm{Y}_{1}\in\Delta_{1},\mathrm{Y}_{2}\in\Delta_{2},\text{and }\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset\}$.
+Let $I$ be a nonempty set of indices.
+Define $\bigodot\{\Delta_{i}\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\}$ as expected.
+Note $\bigodot\{\Delta\}=\Delta$.
 Definition 16 (Single-coalition-first neighborhood models).
@@
   • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
-  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST})-\{\emptyset\})$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a general cover of $\mathtt{suc}(s)$.
+  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$.
   • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Definition 17 (Neighborhood functions).
+Definition 17 (Neighborhood functions of single-coalition-first neighborhood models).
 Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
 For every $\mathrm{C}\subseteq\mathtt{AG}$, define the neighborhood function $\mathtt{nei}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
-$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathrm{C}=\emptyset\text{ and }\mathtt{suc}(s)\neq\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathrm{C}\neq\emptyset\end{cases}$
-Fact 6.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
-For all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
+$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}=\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}\neq\emptyset\end{cases}$
+Fact 4.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$.
 Proof.
 Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
-Assume $\mathtt{suc}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is clearly a general cover of $\mathtt{suc}(s)$.
-Assume $\mathrm{C}=\emptyset$ and $\mathtt{suc}(s)\neq\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is clearly a general cover of $\mathtt{suc}(s)$.
-Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
+Assume $\mathtt{suc}(s)=\emptyset$.
+Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is a cover of $\mathtt{suc}(s)$.
+Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$.
+Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is a cover of $\mathtt{suc}(s)$.
+Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$.
+Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
 Let $t\in\mathtt{suc}(s)$.
-Then there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in X_{a_{1}}$, …, there is $X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in X_{a_{n}}$.
-Then $t\in X_{a_{1}}\cap\dots\cap X_{a_{n}}$.
-Note $X_{a_{1}}\cap\dots\cap X_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$.
-Then $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
+We want to show $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$.
+Note $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$ for every $a\in\mathtt{AG}$.
+Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in\mathrm{X}_{a_{1}}$, …, there is $\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in\mathrm{X}_{a_{n}}$.
+Then $t\in\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$.
+Note $\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$.
+Then $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$.
 Definition 18 (Seriality, independence, and determinism of single-coalition-first neighborhood models).
@@
 We say:
-  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if and only if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
-  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if and only if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
-  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if and only if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
-As above, we let the eight strings in the set $\mathrm{ES}$ correspond to the eight combinations of the three properties, respectively.
-For any single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathtt{X}\in\mathrm{ES}$, we say $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model if it has the properties corresponding to $\mathtt{X}$.
+  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
+  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
+  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
+As above, we let the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ in the set $\mathrm{ES}$ correspond to the eight combinations of the three properties, respectively.
+For any $\mathtt{X}\in\mathrm{ES}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, we say $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model if it has the properties corresponding to $\mathtt{X}$.
 Remarks
 Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems.
-Models based on so-called tight alternating transition systems are single-coalition-first neighborhood $\mathtt{SID}$-models.
+Goranko and Jamroga [4] discussed tight alternating transition systems, which we can verify are single-coalition-first neighborhood $\mathtt{SID}$-models.
 4.4 Representation of single-coalition-first action models by single-coalition-first neighborhood models
@@
 Proof.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood $\mathtt{X}$-model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
 Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as follows:
-  • $\mathtt{AC}=\{\alpha_{a-s-X}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{ and }X\in\mathtt{nei}_{a}(s)\}$.
-  • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-X}\in\mathtt{AC}$:
-    $\mathsf{out}_{a}(s,\alpha_{x-u-X})=\begin{cases}X&\text{if $a=x$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
+  • $\mathtt{AC}=\{\alpha_{a-s-\mathrm{X}}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{and }\mathrm{X}\in\mathtt{nei}_{a}(s)\}$.
+  • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-\mathrm{X}}\in\mathtt{AC}$,
+    $\mathsf{out}_{a}(s,\alpha_{x-u-\mathrm{X}})=\begin{cases}\mathrm{X}&\text{if $x=a$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
 Let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
@@
 Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
-It suffices to show that $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
-Assume $\mathrm{C}=\emptyset$.
-Then, $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}=\mathtt{nei}_{\mathrm{C}}(s)$.
-Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
-Note:
-$\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$
-$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{\mathrm{C}}(s,\sigma_{a_{n}})$
-$\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}$
-Let $X\in\mathtt{AE}(s,\mathrm{C})$.
-Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $X=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
-Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, where $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}},\dots,\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$.
-Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}(s,\sigma_{a_{n}})$ and $X$ is not empty.
-Then, $\sigma_{a_{1}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{a_{n}}\in\mathtt{av}_{a_{n}}(s)$.
-Let $1\leq i\leq n$ and $\sigma_{a_{i}}=(a_{i},\alpha_{x-u-X_{i}})$.
-Note $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})\neq\emptyset$.
-Then, $\mathsf{out}(s,\sigma_{a_{i}})=X_{i}$, $\sigma_{a_{i}}=(a_{i},\alpha_{a_{i}-s-X_{i}})$ and $X_{i}\in\mathtt{nei}_{a_{i}}(s)$.
-Then, $X_{1}\cap\dots\cap X_{n}=X$.
-Then, $X\in\mathtt{nei}(s,\mathrm{C})$.
-Let $X\in\mathtt{nei}(s,\mathrm{C})$.
-Then, there are $X_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{n}\in\mathtt{nei}_{a_{n}}(s)$ such that $X_{1}\cap\dots\cap X_{n}=X$.
-Then, $\alpha_{a_{1}-s-X_{1}}\in\mathtt{AC},\dots,\alpha_{a_{n}-s-X_{n}}\in\mathtt{AC}$.
-Let $\sigma_{a_{1}}=\{(a_{1},\alpha_{a_{1}-s-X_{1}})\},\dots,\sigma_{a_{n}}=\{(a_{n},\alpha_{a_{n}-s-X_{n}})\}$.
-Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})=X_{1},\dots,\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})=X_{n}$.
+It suffices to show $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
+Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\text{ and }\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\}$.
+Assume $\mathtt{suc}(s)=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\emptyset$ and $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
+Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$.
+Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$ and $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$.
+Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
+Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
+Let $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$.
+Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $\mathrm{X}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
 Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
-Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=X_{1}\cap\dots\cap X_{n}=X$.
-Note $X$ is not empty.
-Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$.
-Then, $X\in\mathtt{AE}_{\mathrm{C}}(s)$.
+Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{a_{n}})$ and $\mathrm{X}\neq\emptyset$.
+Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\neq\emptyset$, …, $\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\neq\emptyset$.
+Let $i$ be such that $1\leq i\leq n$ and $\sigma_{a_{i}}=\alpha_{x-u-\mathrm{X}_{i}}$.
+By the definition of $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})$, $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})=\mathrm{X}_{i}$, $x=a_{i}$, and $u=s$.
+Then $\sigma_{a_{i}}=\alpha_{a_{i}-s-\mathrm{X}_{i}}$ and $\mathrm{X}_{i}\in\mathtt{nei}_{a_{i}}(s)$.
+Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$.
+Then, $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$.
+Let $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$.
+Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$ for some $\mathrm{X}_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{n}\in\mathtt{nei}_{a_{n}}(s)$.
+Then, $\alpha_{a_{1}-s-\mathrm{X}_{1}}\dots,\alpha_{a_{n}-s-\mathrm{X}_{n}}\in\mathtt{AC}$.
+Note $\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})=\mathrm{X}_{1}$, $\dots$, $\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{n}$.
+Let $\sigma_{\mathrm{C}}=\alpha_{a_{1}-s-\mathrm{X}_{1}}\cup\dots\cup\alpha_{a_{n}-s-\mathrm{X}_{n}}$.
+Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}=\mathrm{X}$.
+Note $\mathrm{X}$ is not empty.
+Then, $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$.
 Theorem 7.
-For every single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, if $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$, then for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
+For every single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$ and single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$, if $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$, then for every $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
 Proof.
@@
 Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
-Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
-First, it is easy to see the following statements are equivalent:
+Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Then, for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
+Let $\mathtt{X}\in\mathrm{ES}$. It is easy to see that the following three groups of statements are equivalent, respectively.
+  1. (1)
     • $\mathtt{S}\textbf{-}\mathtt{AM}$ is serial;
@@
     • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial.
-Second, it is easy to check the following statements are equivalent:
+  2. (2)
     • $\mathtt{S}\textbf{-}\mathtt{AM}$ is independent;
-  • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset,\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
+    • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset,\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
     • for all $s\in\mathtt{ST},\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, and $\mathrm{Y}_{1},\mathrm{Y}_{2}\subseteq\mathtt{ST}$, if $\mathrm{Y}_{1}\in\mathtt{AE}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{AE}_{\mathrm{D}}(s)$, then $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$;
@@
     • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent.
-Third, it is easy to see the following statements are equivalent:
+  3. (3)
     • $\mathtt{S}\textbf{-}\mathtt{AM}$ is deterministic;
@@
     • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic.
-Then, it follows that for all $\mathtt{X}\in\mathrm{ES}$, $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
-Theorem 8 (Representation of single-coalition-first action models by single-coalition-first neighborhood models).
+It follows that $\mathtt{S}\textbf{-}\mathtt{AM}$ is an $\mathtt{X}$-model if and only if $\mathtt{S}\textbf{-}\mathtt{NM}$ is an $\mathtt{X}$-model.
+The following result follows from Theorems 5, 6, and 7.
+Theorem 8 (Representation of the class of single-coalition-first action $\mathtt{X}$-models by the class of single-coalition-first neighborhood $\mathtt{X}$-models).
 For every $\mathtt{X}\in\mathrm{ES}$, the class of single-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of single-coalition-first neighborhood $\mathtt{X}$-models.
-Proof.
-Let $\mathtt{X}\in\mathrm{ES}$.
-Let $\mathtt{S}\textbf{-}\mathtt{AM}$ be a single-coalition-first action $\mathtt{X}$-model.
-By Theorem 5, $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$.
-By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{NM}$ is a single-coalition-first neighborhood $\mathtt{X}$-model.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}$ be a single-coalition-first neighborhood $\mathtt{X}$-model.
-By Theorem 6, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}$.
-By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a single-coalition-first action $\mathtt{X}$-model.
 5 Clear grand-coalition-first action models and clear single-coalition-first neighborhood models
@@
 We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
+The following result offers two alternative definitions of clear grand-coalition-first action models.
 Theorem 9.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, and let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
 The following three conditions are equivalent:
   1. (1)
-    for all $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, if $\sigma_{a}\neq\sigma^{\prime}_{a}$, then $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})=\emptyset$;
+    for all $a\in\mathtt{AG}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, if $\sigma_{a}\neq\sigma^{\prime}_{a}$, then $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})=\emptyset$;
   2. (2)
-    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$;
+    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$;
   3. (3)
-    for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
+    for all $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
 Proof.
@@
 Assume (1).
-Assume (2) does not hold.
-Then, there is $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\neq\emptyset$.
-Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
-Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
-Then, there is $a\in\mathrm{C},\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\subseteq\sigma_{\mathrm{C}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathrm{C}}$ and $\sigma_{a}\neq\sigma^{\prime}_{a}$.
+Suppose (2) does not hold.
+Then, there is $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\neq\sigma^{\prime}_{\mathrm{C}}$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\neq\emptyset$.
+Then, there is $a\in\mathrm{C}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\subseteq\sigma_{\mathrm{C}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathrm{C}}$, and $\sigma_{a}\neq\sigma^{\prime}_{a}$.
+Let $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$.
+Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{\mathrm{C}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$, and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
 Then, $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
-Then, $w\in\mathsf{out}_{a}(s,\sigma_{a})$ and $w\in\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$.
-Then, $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$.
-There is a contradiction.
+Then, $t\in\mathsf{out}_{a}(s,\sigma_{a})$ and $t\in\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$.
+Then, $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$, which is impossible.
 (2) $\Rightarrow$ (3)
@@
 Assume (3).
-Assume (1) does not hold.
-Then, there is $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\neq\sigma^{\prime}_{a}$ and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$.
-Let $w\in\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$.
-Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
-Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$.
-There is a contradiction.
+Suppose (1) does not hold.
+Then, there is $a\in\mathtt{AG}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\neq\sigma^{\prime}_{a}$ and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$.
+Let $t\in\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$.
+Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$, and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
+Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$, which is impossible.
+For any set $\mathrm{X}$ of states, we say $\Delta\subseteq\mathcal{P}(\mathrm{X})$ is a general partition of $\mathrm{X}$ if (1) $\bigcup\Delta=\mathrm{X}$ and (2) for every $\mathrm{Y},\mathrm{Y}^{\prime}\in\Delta$, if $\mathrm{Y}\neq\mathrm{Y}^{\prime}$, then $\mathrm{Y}\cap\mathrm{Y}^{\prime}=\emptyset$.
+Note a general partition $\Delta$ is a partition if $\emptyset\notin\Delta$.
+Fact 5.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+Then, if $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, then for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$.
+Proof.
+Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear.
+Let $\mathrm{C}\subseteq\mathtt{AG}$ and $t\in\mathtt{ST}$.
+Note $\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$.
+It is easy to show $\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}=\mathtt{suc}(s)$.
+Let $\mathrm{X}_{1},\mathrm{X}_{2}\in\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ such that $\mathrm{X}_{1}\neq\mathrm{X}_{2}$.
+Then, $\mathrm{X}_{1}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{1})$ for some $\sigma_{\mathrm{C}}^{1}\in\mathtt{JA}_{\mathrm{C}}$ and $\mathrm{X}_{2}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{2})$ for some $\sigma_{\mathrm{C}}^{2}\in\mathtt{JA}_{\mathrm{C}}$.
+Note $\sigma^{1}_{\mathrm{C}}\neq\sigma^{2}_{\mathrm{C}}$.
+Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$, that is, $\mathrm{X}_{1}\cap\mathrm{X}_{2}=\emptyset$.
+Note the other direction of the statement of this fact might not hold.
+Suppose $\mathtt{AG}=\{a\}$.
+Consider a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ such that:
+  • $\mathtt{ST}=\{s\}$;
+  • $\mathtt{AC}=\{\alpha_{1},\alpha_{2}\}$;
+  • $\mathsf{out}_{a}(s,\alpha_{1})=\mathsf{out}_{a}(s,\alpha_{2})=\mathtt{ST}$.
+It is easy to check for all $\mathrm{C}\subseteq\mathtt{AG}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$.
+However, $\mathtt{G}\textbf{-}\mathtt{AM}$ is not clear.
 Theorem 10.
@@
 Proof.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
-To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is essentially a single-coalition-first action model, we need to show:
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model, we need to show:
   1. (1)
-    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
+    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
   2. (2)
-    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
+    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
     $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
@@
     $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
-By Fact 5, (1) holds. Clearly, (3) holds. It remains to show (2).
-Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
-Assume $\mathrm{C}=\emptyset$. By the definition of $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathtt{suc}(s)$;
-Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
-By Fact 4, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
-Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
+By Fact 2, (1) holds. Clearly, (3) holds. It remains to show (2). Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
+Assume $\mathrm{C}=\emptyset$.
+Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{\mathtt{AG}}\}=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}=\mathtt{suc}(s)$;
+Assume $\mathrm{C}\neq\emptyset$.
+Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
+By Fact 3, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
+Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
 Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$.
@@
 Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
-It is easy to find a single-coalition-first action model with a $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, $\sigma_{a}\neq\sigma^{\prime}_{a}$, and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$.
-Then it is not a clear grand-coalition-first action model.
+It is easy to find a single-coalition-first action model that is not clear.
 Remarks
 Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models.
-As observed by Ågotnes and Alechina, the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
+As observed by Ågotnes and Alechina [12], the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
 5.2 Clear single-coalition-first neighborhood models
@@
 Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
-We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$, and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
+We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
+The following result gives an alternative definition of clear single-coalition-first neighborhood models.
 Theorem 11.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
 The following two conditions are equivalent:
   1. (1)
-    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
+    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
   2. (2)
-    for all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
+    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
 Proof.
 (1) $\Rightarrow$ (2)
-Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
-By Fact 6, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$.
+Assume (1).
+Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
+By Fact 4, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$.
 Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$.
 Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
-Without loss of any generality, let $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$.
+Without loss of any generality, we can assume there is $w$ such that $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$.
 Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$.
-Note $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$. Then $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
-Then, there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=X_{a_{1}}\cap\dots\cap X_{a_{n}}$, and there is $X^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=X^{\prime}_{a_{1}}\cap\dots\cap X^{\prime}_{a_{n}}$.
-Then $w\notin X^{\prime}_{a_{i}}$ for some $1\leq i\leq n$.
-Note $w\in X_{a_{i}}$.
-Then $X_{a_{i}}\neq X^{\prime}_{a_{i}}$.
-Note $u\in X_{a_{i}}$ and $u\in X^{\prime}_{a_{i}}$.
-Then $X_{a_{i}}\cap X^{\prime}_{a_{i}}\neq\emptyset$.
+It is easy to see $\mathtt{suc}(s)\neq\emptyset$.
+Then, $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$, which is a singleton.
+Note $\mathtt{nei}_{\mathrm{C}}(s)$ is not a singleton.
+Then, $\mathrm{C}\neq\emptyset$.
+Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
+Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$, and there is $\mathrm{X}^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=\mathrm{X}^{\prime}_{a_{1}}\cap\dots\cap\mathrm{X}^{\prime}_{a_{n}}$.
+Note $w\notin\mathrm{Y}_{2}$.
+Then $w\notin\mathrm{X}^{\prime}_{a_{i}}$ for some $a_{i}$.
+Note $w\in\mathrm{X}_{a_{i}}$.
+Then $\mathrm{X}_{a_{i}}\neq\mathrm{X}^{\prime}_{a_{i}}$.
+Note $u\in\mathrm{X}_{a_{i}}$ and $u\in\mathrm{X}^{\prime}_{a_{i}}$.
+Then $\mathrm{X}_{a_{i}}\cap\mathrm{X}^{\prime}_{a_{i}}\neq\emptyset$.
 We have a contradiction.
@@
 (1) is a special case of (2).
-By 9, clear grand-coalition-first action models can be equivalently defined by three different ways, which respectively concern all agents, all coalitions and the grand coalition.
-It would be expected that the following condition is equivalent to the three conditions given in Theorem 11: for all $\mathtt{AG}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$.
-This is actually not the case.
-What follows is a counter-example.
+By Theorem 9, clear grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition.
+It might be expected that the following condition, which concerns the grand coalition, is equivalent to the two conditions given in Theorem 11: for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$.
+Actually, this is not the case.
+What follows is a counterexample.
 Assume $\mathtt{AG}=\{a,b\}$.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, where:
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model such that:
   • $\mathtt{ST}=\{s_{0},s_{1},s_{2},s_{3}\}$;
@@
 Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems.
-It can be shown that those models are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
+It can be shown that they are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
 5.3 Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models
@@
 By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model.
-It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model.
+It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model.
 ∎
 Theorem 13.
-Every clear grand-coalition-first neighborhood model $z$-represents a clear single-coalition-first action model.
+Every clear single-coalition-first neighborhood model $z$-represents a clear grand-coalition-first action model.
 Proof.
@@
 As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
-It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
-From Theorem 12, Theorem 13, and Theorem 7, the following result follows:
-Theorem 14 (Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models).
+It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear.
+By Theorem 4, it is a grand-coalition-first action model.
+Theorem 14 (Representation of the class of clear grand-coalition-first action $\mathtt{X}$-models by the class of clear single-coalition-first neighborhood $\mathtt{X}$-models).
 For every $\mathtt{X}\in\mathrm{ES}$, the class of clear grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of clear single-coalition-first neighborhood $\mathtt{X}$-models.
-7 Tree-like grand-coalition-first action models and tree-like single-coalition-first neigh-borhood models
+Proof.
+Let $\mathtt{X}\in\mathrm{ES}$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action $\mathtt{X}$-model.
+By Theorem 12, $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$.
+By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model.
+By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood $\mathtt{X}$-model.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}$ be a clear single-coalition-first neighborhood $\mathtt{X}$-model.
+By Theorem 13, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents a clear grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$.
+By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model.
+By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a clear grand-coalition-first action $\mathtt{X}$-model.
+6 Tree-like grand-coalition-first action models and tree-like single-coalition-first neighborhood models
 6.1 Tree-like grand-coalition-first action models
 Definition 21 (Histories in grand-coalition-first action models).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
-For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
-Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
+Specially, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
 Definition 22 (Tree-like grand-coalition-first action models).
 Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
-We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history $\theta$ from $r$ to $s$.
+We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
+Note every tree-like model has a unique root.
+The following result gives two alternative definitions of tree-like grand-coalition-first action models.
 Theorem 15.
@@
   1. (1)
-    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$;
+    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
   2. (2)
-    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
+    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
   3. (3)
-    there is a state $r$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
+    there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
 Proof.
@@
 Assume (1).
-Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $a$-history from $r$ to $s$.
-Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $a\in\mathtt{AG}$.
-  • Assume $s=r$.
-Then, $r$ is an $\mathtt{AG}$-history from $r$ to $r$.
-Assume $(r,\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{m},r)$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$.
-Then, $(r,\sigma_{\mathtt{AG}}^{1}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{m}|_{a},r)$ is an $a$-history from $r$ to $r$ which is different from $r$.
+Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$.
+Let $s\in\mathtt{ST}$.
+We want to show that there is an unique $\mathtt{AG}$-history from $r$ to $s$.
+Let $a\in\mathtt{AG}$.
+Assume $s=r$. Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$.
+Assume there is an $\mathtt{AG}$-history $\theta_{\mathtt{AG}}$ from $r$ to $r$ which is different from $r$.
+It is impossible that $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$.
+Assume $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$.
+It is easy to check that $\theta_{a}=(s_{0},\sigma_{\mathtt{AG}}^{1}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ is an $a$-history from $r$ to $r$.
+Clearly, $\theta_{a}$ is different from $r$.
 Note $r$ is an $a$-history from $r$ to $r$.
@@
 Assume $s\neq r$.
-Let $(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$ be the $a$-history from $r$ to $s$.
-For every $1\leq i\leq n$, let $\sigma^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{a}\subseteq\sigma^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{i}_{\mathtt{AG}})$.
-Then $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $s$.
-    Let $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ be an $\mathtt{AG}$-history from $r$ to $s$.
-We claim $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},$ $s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are identical.
-    Assume $n\neq m$.
-Let $a\in\mathtt{AG}$.
-Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$.
-Then they are different.
-We have a contradiction.
+Let $\theta_{a}$ be the unique $a$-history from $r$ to $s$.
+Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
+Assume $\theta_{a}=(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$.
+For every $i$ such that $1\leq i\leq n$, let $\sigma^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{a}\subseteq\sigma^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\sigma^{i}_{\mathtt{AG}})$.
+Then, $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $s$.
+Assume there is a different $\mathtt{AG}$-history $\theta^{\prime}_{\mathtt{AG}}$ from $r$ to $s$.
+Note it is impossible $\theta^{\prime}_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta^{\prime}_{\mathtt{AG}}=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
+We want to show $(s_{0},\sigma^{1}_{\mathtt{AG}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
+Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$.
+Then $n=m$ and $s_{i}=t_{i}$ for all $i$.
 Assume $\sigma^{i}_{\mathtt{AG}}\neq\lambda^{i}_{\mathtt{AG}}$ for some $1\leq i\leq n$.
-Then there is $a\in\mathtt{AG}$ such that $\sigma^{i}_{\mathtt{AG}}|_{a}\neq\lambda^{i}_{\mathtt{AG}}|_{a}$.
-Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(s_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$.
-Then, they are different.
-We have a contradiction.
-    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$.
-Let $a\in\mathtt{AG}$.
-Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{a},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{a},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{a},t_{1},\dots,$ $\lambda_{\mathtt{AG}}^{m}|_{a},t_{m})$ are $a$-histories from $r$ to $s$.
-Then, they are different.
+Then, there is $b\in\mathtt{AG}$ such that $\sigma^{i}_{\mathtt{AG}}|_{b}\neq\lambda^{i}_{\mathtt{AG}}|_{b}$.
+Note both $(s_{0},\sigma^{1}_{\mathtt{AG}}|_{b},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{b},s_{n})$ and $(t_{0},\lambda^{1}_{\mathtt{AG}}|_{b},t_{1},\dots,\lambda_{\mathtt{AG}}^{m}|_{b},t_{m})$ are $b$-histories from $r$ to $s$, which are different.
 We have a contradiction.
 (3) $\Rightarrow$ (2)
-Assume (3). Let $r$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ such that for every state $s$ of $\mathtt{G}\textbf{-}\mathtt{AM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
-Let $s$ be a state of $\mathtt{G}\textbf{-}\mathtt{AM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
-  • Assume $s=r$.
-Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$.
-Assume $(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$.
-For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\lambda^{i}_{\mathtt{AG}})$.
-Then $(s_{0},\lambda^{1}_{\mathtt{AG}},s_{1},\dots,\lambda_{\mathtt{AG}}^{m},s_{m})$ is an $\mathtt{AG}$-history from $r$ to $r$ which is different from $r$.
+Assume (3).
+Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
+Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
+We want to show that there is an unique $\mathrm{C}$-history from $r$ to $s$.
+Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$.
+Assume there is a $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$ which is different from $r$.
+Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Assume $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$.
+For every $1\leq i\leq n$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $s_{i}\in\mathsf{out}_{\mathtt{AG}}(s_{i-1},\lambda^{i}_{\mathtt{AG}})$.
+Then, $(s_{0},\lambda^{1}_{\mathtt{AG}},s_{1},\dots,\lambda_{\mathtt{AG}}^{n},s_{n})$ is an $\mathtt{AG}$-history from $r$ to $r$, which is different from $r$.
 Note $r$ is an $\mathtt{AG}$-history from $r$ to $r$.
@@
 Assume $s\neq r$.
-Let $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$ be the $\mathtt{AG}$-history from $r$ to $s$.
-Then, $,(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},$$\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is an $\mathrm{C}$-history from $r$ to $s$.
-    Assume $(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$ which is different from $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},$ $s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$.
-    For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$.
-Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
-    Then $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is different from $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$.
+Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$.
+Note it is impossible $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$.
+It is easy to see $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
+Assume there is a different $\mathrm{C}$-history $\theta^{\prime}_{\mathrm{C}}$ from $r$ to $s$.
+Again, it is impossible $\theta_{\mathrm{C}}^{\prime}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta^{\prime}_{\mathrm{C}}=(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$.
+For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$.
+Then, $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
+It is easy to check that $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is different from $(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$.
 We have a contradiction.
 (2) $\Rightarrow$ (1)
-It is easy to see that (1) is a special case of (2).
+(1) is a special case of (2).
 Theorem 16.
@@
 Proof.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model.
-We are done if we can show for all $s\in\mathtt{ST}$, and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
-Let $s\in\mathtt{ST}$, $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model and $r$ be its root.
+Let $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$.
+It suffices to show $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
 Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$.
-Let $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
-Then, there are at least two paths $(s,\sigma_{\mathtt{AG}},w)$ and $(s,\sigma^{\prime}_{\mathtt{AG}},w)$ from $s$ to $w$.
-Then, there are at least two paths $(r,\dots,s,\sigma_{\mathtt{AG}},w)$ and $(r,\dots,s,\sigma^{\prime}_{\mathtt{AG}},w)$ from r to w.
-Note that $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like grand-coalition-first action model.
-Then, there is a contradiction.
-It is easy to find a clear grand-coalition-first action model with a loop. Then, it is not tree-like.
+Let $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$.
+Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$.
+It is easy to check that both $(\theta_{\mathtt{AG}},\sigma_{\mathtt{AG}},t)$ and $(\theta_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}},t)$ are $\mathtt{AG}$-histories from $r$ to $t$, which are different.
+We have a contradiction.
+It is easy to find a clear grand-coalition-first action model with a loop, which is not tree-like.
 Remarks
-In the literature, Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models.
-They are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined here.
-It is required in their definition that for every state, there is a unique computation, which is a sequence of states, from the root to the state.
-Consequently, they are not clear.
+Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models, which are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined in this section.
+For the former, it is only required that for every state, there is a unique computation, which is a sequence of states, from the root to the state.
 6.2 Tree-like single-coalition-first neighborhood models
 Definition 23 (Histories in single-coalition-first neighborhood models).
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
-For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
 Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
@@
 Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
-We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ from $r$ to $s$.
+We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
+The following result is easy to verify and we skip its proof.
+Fact 6.
+Every tree-like single-coalition-first neighborhood model is a clear single-coalition-first neighborhood model, but not vice versa.
+The following result offers an alternative definition of tree-like single-coalition-first neighborhood models.
 Theorem 17.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
 The following two conditions are equivalent:
   1. (1)
-    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$;
+    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
   2. (2)
-    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
+    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
 Proof.
 (1) $\Rightarrow$ (2)
-Assume (1). Let $r$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$.
-Let $s$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
+Assume (1).
+Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
+Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
+We want to show that there is a unique $\mathrm{C}$-history from $r$ to $s$.
 Assume $s=r$.
-Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$.
-Assume $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$.
-For every $1\leq i\leq m$, let $Z_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $Z_{i}\subseteq\mathrm{Y}_{i}$ and $s_{1}\in Z_{i}$.
-Then, $(s_{0},Z_{1},s_{1},\dots,Z_{m},s_{m})$ is an $a$-history from $r$ to $r$ which is different from $r$.
+Note $r$ is a $\mathrm{C}$-history from $r$ to $r$.
+Assume there is a different $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$.
+Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
+Assume $\mathrm{C}=\emptyset$.
+Then for every $i$ such that $1\leq i\leq n$, $\mathrm{Y}_{i}=\mathtt{suc}(s_{i-1})$.
+Let $a\in\mathtt{AG}$.
+For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$.
+Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$.
+Note $r$ is an $a$-history from $r$ to $r$.
+We have a contradiction.
+Assume $\mathrm{C}\neq\emptyset$.
+Let $a\in\mathrm{C}$.
+Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$ for every $s\in\mathtt{ST}$.
+For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq\mathrm{Z}_{i}$.
+Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{m},s_{m})$ is an $a$-history from $r$ to $r$, which is different from $r$.
 Note $r$ is an $a$-history from $r$ to $r$.
@@
 Assume $\mathrm{C}=\emptyset$.
-Let $a\in\mathtt{AG}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$.
-It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$ is a unique $a$-history from $r$ to $s$.
-    Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$.
-    For every $1\leq i\leq n$, let $Z_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq Z_{i}$.
-Then, $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
-    Let $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ be a $\mathrm{C}$-history from $r$ to $s$. We claim $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ and $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ are identical.
-    For every $1\leq i\leq m$, let $E_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $D_{i}\subseteq E_{i}$. Then, $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$ is a $a$-history from $r$ to $s$.
-    Assume $n\neq m$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
-    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$.
-Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$.
-We have a contradiction.
-    Assume $Z_{i}\neq D_{i}$ for some $1\leq i\leq n$.
-Then, $Z_{i}\cap D_{i}=\emptyset$.
-Then, $s_{i}\neq t_{i}$.
-Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,$ $\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$.
-We have a contradiction.
+Let $a\in\mathtt{AG}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$.
+It is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
+It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),$ $s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
+Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$.
+We want to show $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
+It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta_{\mathrm{C}}=(t_{0},\mathtt{suc}(t_{0}),t_{1},\dots,\mathtt{suc}(t_{m-1}),t_{m})$.
+For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathtt{suc}(t_{i-1})$ and $t_{i}\in\mathrm{Z}_{i}$.
+Then, $(t_{0},\mathrm{Z}_{1},t_{1},\dots,\mathrm{Z}_{m},t_{m})$ is an $a$-history from $r$ to $s$.
+Then $n=m$ and $s_{i}=t_{i}$ for all $0\leq i\leq n$.
+Then, $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
+Assume $\mathrm{C}\neq\emptyset$.
+Let $a\in\mathrm{C}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$.
+Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
+For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$.
+Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
+Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$.
+It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta_{\mathrm{C}}=(t_{0},\mathrm{U}_{1},t_{1},\dots,\mathrm{U}_{m},t_{m})$.
+For every $1\leq i\leq m$, let $\mathrm{V}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{U}_{i}\subseteq\mathrm{V}_{i}$.
+Then, $(t_{0},\mathrm{V}_{1},t_{1},\dots,\mathrm{V}_{m},t_{m})$ is an $a$-history from $r$ to $s$.
+Then $n=m$ and $s_{i}=t_{i}$ for every $0\leq i\leq n$.
+Assume $\mathrm{Z}_{i}\neq\mathrm{U}_{i}$ for some $1\leq i\leq n$.
+Note $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood model.
+Then, $\mathrm{Z}_{i}\cap\mathrm{U}_{i}=\emptyset$.
+Then, $s_{i}\neq t_{i}$, which is impossible.
+Then, $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$.
 (2) $\Rightarrow$ (1)
-It easy to see that (1) is a special case of (2).
-By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which are respectively concerned with all agents, all coalitions, and the grand coalition.
-It would be expected that the following condition is equivalent to the two conditions in Theorem 17: there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathtt{AG}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
+(1) is a special case of (2).
+By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition.
+It might be expected that the following condition is equivalent to the two conditions in Theorem 17: there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
 In fact, this is not the case.
-The counterexample given after Theorem 11 also works here.
+The example given after Theorem 11 also works here.
 6.3 Representation of tree-like grand-coalition-first action models by tee-like single-coalition-first neighborhood models
@@
 Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model.
-By Theorem 4, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model.
-Then, it is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model.
+By Theorems 16 and 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model.
+It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model.
 ∎
 Theorem 19.
-Every tree-like grand-coalition-first neighborhood model $z$-represents a tree-like single-coalition-first action model.
+Every tree-like single-coalition-first neighborhood model $z$-represents a tree-like grand-coalition-first action model.
 Proof.
@@
 Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6.
-As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
-It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.
-From Theorem 18, Theorem 19, and Theorem 7, the following result follows:
-Theorem 20 (Representation of tree-like grand-coalition-first action models by tree-like single-coalition-first neighborhood models).
+As shown there, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
+It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like.
+By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.
+The following result follows from Theorems 18, 19, and 7.
+Theorem 20 (Representation of the class of tree-like grand-coalition-first action $\mathtt{X}$-models by the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models).
 For every $\mathtt{X}\in\mathrm{ES}$, the class of tree-like grand-coalition-first action $\mathtt{X}$-models is $z$-representable by the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models.
 7 Each of those eight coalition logics is determined by these six kinds of models, too
-For every finite nonempty sequence $\chi$, we use $\mathsf{last}(\chi)$ to denote the last element of $\chi$.
 Theorem 21.
-Let $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model.
-There is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ such that for all $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$.
+For every $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, there is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime})$ such that for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime}\Vdash\phi$.
 Proof.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
+Lert $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
+In the following, for every finite nonempty sequence $\pi$, we use $\pi^{l}$ to denote the last element of $\pi$.
 Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},$ $\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ as follows:
   • $\mathtt{ST}^{\prime}=\mathtt{ST}_{0}^{\prime}\cup\mathtt{ST}_{1}^{\prime}\cup\mathtt{ST}_{2}^{\prime}\cup\dots$, where:
-    • $\mathtt{ST}_{0}^{\prime}=\{w\}$;
-    • $\mathtt{ST}_{k+1}^{\prime}=\{\chi-\sigma_{\mathtt{AG}}-u\mid\chi\in\mathtt{ST}_{k}^{\prime},\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}},u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
-  • for every $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$:
-    $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})=\{y\in\mathtt{ST}^{\prime}\mid y=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
-  • for every $\chi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\chi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\mathsf{last}(\chi))\}$.
-Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
-We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is what we want.
+    • $\mathtt{ST}_{0}^{\prime}=\{s\}$;
+    • $\mathtt{ST}_{k+1}^{\prime}=\{\pi-\sigma_{\mathtt{AG}}-u\mid\pi\in\mathtt{ST}_{k}^{\prime},\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}},\text{and }u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\}$.
+  • for every $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})=\{\pi-\sigma_{\mathtt{AG}}-u\mid u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\}$.
+  • for every $\pi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\pi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\pi^{l})\}$.
+Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$ and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
+We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is what we want.
+We first make two claims:
+  1. (1)
+    for every $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$, if $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, then $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$;
+  2. (2)
+    for every $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, then there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
+First, we show the claim (1).
+Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$.
+Assume $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
+Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
+Then, $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$.
+Note $\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\subseteq\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
+Then, $u\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, that is, $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
+Second, we show the claim (2).
+Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $t\in\mathtt{ST}$.
+Assume $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
+Then, $t\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
+Then, there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$.
+Note $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})\subseteq\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
+Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
 We show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is an $\mathtt{X}$-model.
@@
 Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent.
-Let $\chi\in\mathtt{ST}^{\prime}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ and $\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{D}}(\chi)$.
-Then, $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}^{\prime}_{\mathrm{D}}(\chi,\sigma_{\mathrm{D}})\neq\emptyset$.
-    Then, there is $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$.
-Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
-    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$.
-Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$.
-Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$.
-Similarly, we can show $\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(\mathsf{last}(\chi))$.
-Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi))$.
-Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$.
-Let $u^{\prime}\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
-    Then $u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma^{\prime}_{\mathtt{AG}})$ for some $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
-    Then $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma^{\prime}_{\mathtt{AG}})$.
-    Then, $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
-Then, $\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$.
-Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\chi)$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is independent.
-  • Assume $(\mathtt{G}\textbf{-}\mathtt{AM},w)$ is deterministic.
+Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$, and $\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{D}}(\pi)$.
+To show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is independent, it suffices to show $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi)$.
+Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}^{\prime}_{\mathrm{D}}(\pi,\sigma_{\mathrm{D}})\neq\emptyset$.
+By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}_{\mathrm{D}}(\pi^{l},\sigma_{\mathrm{D}})\neq\emptyset$.
+Then $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ and $\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(\pi^{l})$.
+Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\pi^{l})$.
+Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\pi^{l},\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$.
+By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$.
+Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi)$.
+Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic.
 Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic.
-Then, there is $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\chi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ is not a singleton.
-Let $s,s^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ where $s\neq s^{\prime}$.
-Let $s=\chi-\sigma_{\mathtt{AG}}-u$ and $s^{\prime}=\chi-\sigma_{\mathtt{AG}}-u^{\prime}$ where $u\neq u^{\prime}$.
-Then, $u,u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$.
-Then, there is a contradiction.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is deterministic.
-It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is a tree-like model.
-We show for all $\chi$, and $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},w\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w\Vdash\phi$.
-We put an induction on $\phi$. We consider only the cases $\phi=p$ and $\phi=\langle[\mathrm{C}]\rangle\psi$, and skip other cases.
-  • $\phi$ is a proposition letter $p$.
-    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash p$.
-Then, $p\in\mathsf{label}^{\prime}(\chi)$.
-Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$.
-    $\Rightarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash p$.
-Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$.
-Then, $p\in\mathsf{label}^{\prime}(\chi)$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$.
-  • $\phi$ is $\langle[\mathrm{C}]\rangle\psi$.
-    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\langle[\mathrm{C}]\rangle\psi$.
-Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ such that for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
-    We are done if we can show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
-    Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$.
-Let $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$.
-Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
-    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$.
-Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$.
-Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$.
-    Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$.
-Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\text{and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
-    Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$.
-    Then, $y\in\mathsf{out}^{\prime}(\chi,\sigma_{\mathrm{C}})$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
-By the inductive hyphothesis, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(y)\Vdash\psi$.
-That is $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
-    $\Rightarrow$
-    Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\langle[\mathrm{C}]\rangle\psi$.
-Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
-    We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
-    Note $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$.
-Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$.
-Then $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
-Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$.
-Then $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$.
-Then $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$ is not empty.
-Then $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$.
-    Let $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$.
-Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})\text{ and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
-    Then, there is $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ such that $t=\mathsf{last}(y)$.
-    Then, $t\in\mathsf{out}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$.
-Then, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
-By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
+Then, there is $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\pi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ is not a singleton.
+Let $\pi^{\prime},\pi^{\prime\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ such that $\pi^{\prime}\neq\pi^{\prime\prime}$.
+Let $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime}$ and $\pi^{\prime\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime\prime}$.
+Note $t^{\prime}\neq t^{\prime\prime}$.
+Then, $t^{\prime},t^{\prime\prime}\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$.
+We have a contradiction.
+It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is tree-like.
+We show for all $\pi\in\mathtt{ST}^{\prime}$ and $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\phi$, which implies for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$.
+We put an induction on $\phi$.
+We consider only the case $\phi=\langle[\mathrm{C}]\rangle\psi$ and skip others.
+Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\langle[\mathrm{C}]\rangle\psi$.
+Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ such that for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
+We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
+Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$.
+By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$.
+Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$.
+Let $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
+By the claim (2), there is a $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
+Note $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
+By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
+Assume $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\langle[\mathrm{C}]\rangle\psi$.
+Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
+We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ and for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
+Note $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$.
+By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$.
+Then $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$.
+Let $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
+By the claim (1), $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
+Note $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{\prime l}\Vdash\psi$.
+By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
 Theorem 22.
-For every $\mathtt{X}\in\mathrm{ES}$ and $\phi\in\Phi$, the following statements are equivalent:
+For all $\mathtt{X}\in\mathrm{ES}$ and $\phi\in\Phi$, the following statements are equivalent:
   1. (0)
@@
 Proof.
-By Theorem 8 and Theorem 1, (2), and (3) are equivalent.
-By Theorem 14 and Theorem 1, (4), and (5) are equivalent.
-By Theorem 20 and Theorem 1, (6), and (7) are equivalent.
+By Theorem 8 and Theorem 1, (1) and (2) are equivalent.
+By Theorem 14 and Theorem 1, (3) and (4) are equivalent.
+By Theorem 20 and Theorem 1, (5) and (6) are equivalent.
 It suffices to show that (0), (1), (3) and (5) are equivalent.
@@
 By Theorem 16, (3) implies (5).
-Suppose (6).
-Assume (1) does not hold.
-Then $\phi$ is false at a pointed grand-coalition-first action model $(\mathtt{G}\textbf{-}\mathtt{AM},w)$.
-Then, by Theorem 21, there is a tree-like grand-coalition-first model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ such that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)\nVdash\phi$.
-We have a contradiction.
-Then, (1).
-9 Concluding remarks
-Grand-coalition-first action models are those action models in which the outcome function for the grand coalition determines the outcome functions for other coalitions.
-Single-coalition-first action models are those action models in which the outcome functions for single coalitions determine the outcome functions for other coalitions.
-This class of models can be $z$-represented by the class of single-coalition-first neighborhood models.
-Clear grand-coalition-first action models are those single-coalition-first action models where different joint actions of a coalition have different outcomes.
-This class of models can be $z$-represented by the class of clear single-coalition-first neighborhood models.
-Tree-like single-coalition-first action models are those single-coalition-first action models where every state has a determined history.
-This class of models can be $z$-represented by the class of tree-like single-coalition-first neighborhood models.
-These kinds of models have different motivations.
-Every kind of model has eight subclasses, determined by which of the properties seriality, independence of agents, and determinism hold.
-In this paper, we show that the subclasses of models that meet the same properties determine the same coalition logic.
+By Theorem 21, (5) implies (1).
+8 Conclusion
+Li and Ju [9, 8] proposed eight coalition logics based on grand-coalition-first action models.
+In this work, we show that each of them is determined by six kinds of models, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
 Acknowledgments

### v3

@@
-Each of those eight coalition logics is also determined by six other kinds of models
+Seven kinds of equivalent models
 Abstract
-Coalition Logic is a central logic in logical research on strategic reasoning.
+Coalition Logic is an important logic in logical research on strategic reasoning.
 In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism.
 They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions are met.
-In this paper, we show that each of the eight logics is also determined by six other kinds of models, with respective properties, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
+In this paper, we show that each of the eight sets of valid formulas of the eight logics is determined by six other kinds of models, that is, single-coalition-first action models, single-coalition-first actual neighborhood models, clear grand-coalition-first action models, clear single-coalition-first actual neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first actual neighborhood models.
 Keywords: coalition logics, action models, neighborhood models, seriality, independence of agents, determinism
 1 Introduction
-1.1 Eight coalition logics determined by eight classes of general concurrent game models
-Coalition Logic $\mathsf{CL}$, proposed by Pauly ([10, 11]), is a central logic in logical research on strategic reasoning.
-The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{C}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{C}$ ensures $\phi$.
-$\mathsf{CL}$ has two kinds of models.
-The first kind is concurrent game models.
+1.1 Coalition Logic and its two kinds of semantics
+Coalition Logic $\mathsf{CL}$, proposed by Pauly [pauly_modal_2002], is an important logic in logical research on strategic reasoning.
+Many logics in this field are extensions of $\mathsf{CL}$.
+For example, Alternating-time Temporal Logic $\mathsf{ATL}$ [alur_alternating-time_2002] is a temporal extension of $\mathsf{CL}$, and Strategy Logic $\mathsf{SL}$ [mogavero_reasoning_2014] is an extension of $\mathsf{CL}$, whose language has quantifiers for and names of strategies.
+We refer to [benthem_models_2015], [agotnes_knowledge_2015] and [sep-logic-power-games] for overviews of the area.
+The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$, indicating some available joint action of coalition $\mathrm{C}$ ensures $\phi$.
+$\mathsf{CL}$ has two kinds of equivalent semantics [pauly_modal_2002, goranko_strategic_2013]: action semantics and neighborhood semantics.
+In action semantics, models are concurrent game models.
 Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; for every coalition, there is an availability function, specifying available joint actions of the coalition at states; for every coalition, there is an outcome function, specifying possible outcome states of joint actions of the coalition.
-The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
-The second kind is alpha neighborhood models.
+The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
+In neighborhood semantics, models are alpha neighborhood models.
 Alpha neighborhood models do not have actions.
@@
 Instead, for every coalition, there is a neighborhood function that specifies the alpha powers of the coalition.
-An alpha power is a set of states into which the coalition can force the next moment to fall.
-The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
-[9] argued that concurrent game models have three too strong assumptions: seriality, the independence of agents, and determinism.
-Based on general concurrent game models, which do not have the three assumptions, they proposed a Minimal Coalition Logic $\mathsf{MCL}$.
-Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some kinds of situations.
+An alpha power is a set of states into which the coalition can force the next moment to fall [moulin_cores_1982].
+The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
+Concurrent game models and alpha neighborhood models of $\mathsf{CL}$ have their respective advantages and disadvantages.
+Concurrent game models are more intuitive.
+However, they contain information for which the language of $\mathsf{CL}$ is blind, that is, actions.
+Neighborhood models are simpler: they only keep the needed information for evaluating formulas of $\mathsf{CL}$.
+Some work in the line of $\mathsf{CL}$, such as [agotnes_coalition_2018], uses neighborhood models
+1.2 Eight coalition logics determined by eight classes of general concurrent game models
+Concurrent game models have three assumptions, among others.
+The first one is seriality: every coalition always has an available joint action.
+The second one is the independence of agents: the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions.
+The third one is determinism: every available joint action of the grand coalition has a unique outcome.
+The three assumptions are also reflected in the alpha neighborhood models of $\mathsf{CL}$.
+In the agency literature, some authors have used models in which one or more of the three assumptions do not hold.
+In the tradition of coalition logic, Jiang and Naumov [JIANG2022103727] used non-deterministic concurrent game models.
+In the tradition of STIT (see to it that) logic, Sergot [sergot_examples_2014, sergot_actual_2022] argued that neither the independence of agents nor determinism commonly holds.
+At the intersection of $\mathsf{ATL}$ and STIT logic, Boudou and Lorini [10.5555/3237383.3237443] utilized non-deterministic concurrent game models.
+In the tradition of Propositional Dynamic Logic, Royakkers and Hughes [royakkers_blame_2020] argued that the independence of agents does not hold generally and used models in which the independence of agents and determinism fail.
+Against the agency literature, Li and Ju [li_minimal_2025] argued that the three assumptions are too strong.
+Here, we briefly mention their arguments.
+Usually, games terminate when reaching specific states.
+For example, a paper-scissors-rock game terminates when a winner is decided.
+Intuitively, in terminating states, players have no available actions.
+Thus, seriality might not hold.
+There are many situations in which whether a coalition can perform an action is conditional on other agents’ actions at the same time.
+Here is an example.
+There are two agents in a room, $a$ and $b$, and only one chair.
+Agent $a$ can sit, and agent $b$ can sit, but they cannot sit at the same time.
+The independence of agents fails in these situations.
+In many situations, some joint actions of all behaving agents have more than one outcome state.
+The following example, from [sergot_examples_2014], can illustrate this.
+A vase stands on a table.
+There is an agent $a$ who can lift or lower the end of the table.
+If the table tilts, the vase might fall, and if it falls, it might break.
+Therefore, determinism might not hold.
+Based on general concurrent game models, which do not have the three assumptions, Li and Ju [li_minimal_2025] proposed a Minimal Coalition Logic $\mathsf{MCL}$.
+Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some situations.
 Considering which of the three properties we want to keep, there are eight coalition logics in total.
-[8] showed the completeness of the eight logics in a uniform way.
+Li and Ju [li_completeness_2024] showed the completeness of the eight logics in a uniform way.
 1.3 Our work
-In this paper, we call general concurrent game models grand-coalition-first action models.
-In this work, we show that each of the eight coalition logics is also determined by six other kinds of models, with respective properties.
+The eight coalition logics have eight sets of valid formulas.
+In this work, we show the following: for each of the eight sets of valid formulas, (1) it is also determined by three other classes of action models, and (2) it is determined by three other classes of actual neighborhood models.
+These different kinds of models make good sense for representing coalitional powers.
+Our work makes it clear what the logics with respect to these kinds of models are.
+Action models are those like concurrent game models.
+In actual neighborhood models, for every coalition, there is a neighborhood function that specifies actual powers of the coalition.
+An actual power is a set of states such that the coalition has a choice such that (1) the choice forces the next moment to fall into the set, and (2) every state in the set is compatible with the choice [benthem_new_2019].
+To ease understanding the differences among the six kinds of models, in this paper, we call general concurrent game models grand-coalition-first action models.
+The reason is that in general concurrent game models, the outcome functions of all coalitions can be derived from the outcome functions of the grand coalition.
+The six kinds of models are as follows:
+  1. 1.
+    1. (a)
       The first kind is single-coalition-first action models.
-In this kind of models, the outcome functions for single coalitions determine the outcome functions for other coalitions.
+In this kind of models, the outcome functions for single coalitions determine those for other coalitions.
 This kind of models makes good sense for situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other.
-The second kind is single-coalition-first neighborhood models.
-In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition.
-An actual power is a minimal scope into which the coalition can force the next moment to fall.
-In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
+    2. (b)
+      The second kind is single-coalition-first actual neighborhood models. In this kind of models, the neighborhood functions for single coalitions determine those for other coalitions.
+  2. 2.
+    1. (a)
       The third kind is clear grand-coalition-first action models.
@@
 Clear grand-coalition-first action models are quite common in game theory.
-The fourth kind is clear single-coalition-first neighborhood models.
-In this kind of models, the actual powers of an agent are pairwise disjoint.
-The fifth kind is tree-like grand-coalition-first action models.
-In this kind of models, every state has a determined history.
-The sixth kind is tree-like single-coalition-first neighborhood models.
-In this kind of models, every state has a determined history, too.
-The rest of the paper is structured as follows.
+    2. (b)
+      The fourth kind is clear single-coalition-first actual neighborhood models. In this kind of models, the actual powers of an agent are pairwise disjoint.
+  3. 3.
+    1. (a)
+      The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history.
+    2. (b)
+      The sixth kind is tree-like single-coalition-first actual neighborhood models. In this kind of models, every state has a determined history, too.
+This work offers new models for logics like Alternating-time Temporal Logic or STIT logic, which will be our future work.
+Structure of the paper
   • In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from action semantics to neighborhood semantics.
   • In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models.
-In Section 4, we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models and show that they can represent single-coalition-first action models.
-In Section 5, we define clear grand-coalition-first action models and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models and show that they can represent clear grand-coalition-first action models.
-In Section 6, we define tree-like grand-coalition-first action models and show that they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models and show that they can represent tree-like grand-coalition-first action models.
+  • In Section 4, we define single-coalition-first action models and show that they are grand-coalition-first action models.
+    We define single-coalition-first actual neighborhood models and show that they can represent single-coalition-first action models.
+  • In Section 5, we define clear grand-coalition-first action models and show that they are single-coalition-first action models.
+    We define clear single-coalition-first actual neighborhood models and show that they can represent clear grand-coalition-first action models.
+  • In Section 6, we define tree-like grand-coalition-first action models and show that they are clear grand-coalition-first action models.
+    We define tree-like single-coalition-first actual neighborhood models and show that they can represent tree-like grand-coalition-first action models.
   • In Section 7, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties.
@@
 The language $\Phi$ is defined as follows, where $p$ ranges over $\mathtt{AP}$ and $\mathrm{C}\subseteq\mathtt{AG}$:
-$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
-The formula $\langle[\mathrm{C}]\rangle\phi$ indicates some available joint action of $\mathrm{C}$ ensures $\phi$.
-The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual.
+$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$
+The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ indicates some available joint action of $\mathrm{C}$ ensures $\phi$.
+Here, the notation of modalities differs from their notation in the literature.
+We do this to indicate the iteration of quantifiers in their meaning.
+The propositional connectives $\bot,\lor,\rightarrow$ and $\leftrightarrow$ are defined as usual.
 2.2 Action semantics
@@
 Definition 2 (Action models).
-An action model is a tuple $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
+An action model is a tuple $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
   • $\mathtt{ST}$ is a nonempty set of states.
   • $\mathtt{AC}$ is a nonempty set of actions.
-  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an availability function for $\mathrm{C}$.
   • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
+  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an availability function for $\mathrm{C}$ meeting the following condition: for all $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}$.
   • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Different coalition logics might put different constraints on action models.
+Different coalition logics might put various constraints on action models.
 Definition 3 (Action semantics).
+Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
+Then truth conditions of formulas in $\Phi$ at states $s$ are defined recursively as follows:
 $\mathtt{AM},s\Vdash\top$
-$\mathtt{AM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
-$\mathtt{AM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{AM},s\Vdash\phi$
-$\mathtt{AM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{AM},s\Vdash\phi$ and $\mathtt{AM},s\Vdash\psi$
-$\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
+$\mathtt{AM},s\Vdash p$ $\Leftrightarrow$ $p\in\mathsf{label}(s)$
+$\mathtt{AM},s\Vdash\neg\phi$ $\Leftrightarrow$ not $\mathtt{AM},s\Vdash\phi$
+$\mathtt{AM},s\Vdash\phi\land\psi$ $\Leftrightarrow$ $\mathtt{AM},s\Vdash\phi$ and $\mathtt{AM},s\Vdash\psi$
+$\mathtt{AM},s\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ $\Leftrightarrow$ there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
 2.3 Neighborhood semantics
-Definition 4 (Neighborhood models and alpha neighborhood models).
-A neighborhood model is a tuple $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
+Definition 4 (Actual neighborhood models).
+An actual neighborhood model is a tuple $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
   • $\mathtt{ST}$ is a nonempty set of states.
@@
   • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-A neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is closed under supersets.
-Different coalition logics might put different constraints on neighborhood models.
+Intuitively, $\mathtt{nei}_{\mathrm{C}}(s)$ consists of actual powers of $\mathrm{C}$ at $s$.
+Definition 5 (Alpha neighborhood models).
+An alpha neighborhood model is a tuple $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
+  • $\mathtt{ST}$ is a nonempty set of states.
+  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $\mathrm{C}$ such that for every $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is closed under supersets.
+  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
+Intuitively, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ consists of alpha powers of $\mathrm{C}$ at $s$.
+Different coalition logics might put various constraints on neighborhood models.
 Definition 6 (Neighborhood semantics).
+Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an actual/alpha neighborhood model.
+Then truth conditions of formulas in $\Phi$ at states $s$ are defined recursively as follows:
 $\mathtt{NM},s\Vdash\top$
-$\mathtt{NM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
-$\mathtt{NM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{NM},s\Vdash\phi$
-$\mathtt{NM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{NM},s\Vdash\phi$ and $\mathtt{NM},s\Vdash\psi$
-$\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
-The following fact, which is easy to verify, indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing the set of formulas it satisfies.
-Fact 1.
-Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
-Then, for every $s\in\mathtt{ST}$ and $\phi\in\Phi$, $\mathtt{NM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
+$\mathtt{NM},s\Vdash p$ $\Leftrightarrow$ $p\in\mathsf{label}(s)$
+$\mathtt{NM},s\Vdash\neg\phi$ $\Leftrightarrow$ not $\mathtt{NM},s\Vdash\phi$
+$\mathtt{NM},s\Vdash\phi\land\psi$ $\Leftrightarrow$ $\mathtt{NM},s\Vdash\phi$ and $\mathtt{NM},s\Vdash\psi$
+$\mathtt{NM},s\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ $\Leftrightarrow$ there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
 2.4 Transformation of action semantics to neighborhood semantics
 Definition 7 (Actual effectivity functions and alpha effectivity functions of action models).
-Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
+Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
   • For every $\mathrm{C}\subseteq\mathtt{AG}$, the function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ in $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
@@
     $\mathtt{LE}_{\mathrm{C}}(s)=\{\mathrm{Y}\subseteq\mathtt{ST}\mid\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathrm{Y}\text{ for some }\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)\}.$
-Actual effectivity functions are more transparent in the following sense: each element of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions.
-Definition 7 (Action models $z$-representable by neighborhood models and $\alpha$-representable by alpha neighborhood models).
-An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $z$-representable by a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
-A class of action models $\mathbf{AM}$ is $z$-representable by a class of neighborhood models $\mathbf{NM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
-An action model $\mathtt{AM}$ is $\alpha$-representable by an alpha neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{LE}_{\mathrm{C}}=\mathtt{nei}^{\alpha}_{\mathrm{C}}$.
-A class of action models $\mathbf{AM}$ is $\alpha$-representable by a class of alpha neighborhood models $\mathbf{NM}^{\alpha}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}^{\alpha}$, and (2) every alpha neighborhood model in $\mathbf{NM}^{\alpha}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
-Intuitively, that a neighborhood model $\mathtt{NM}$ represents an action model $\mathtt{AM}$ means that $\mathtt{NM}$ contains all the information of $\mathtt{AM}$ to evaluate formulas of $\Phi$.
+Actual effectivity functions are more transparent than alpha effectivity functions in the following sense: every element $\mathrm{Y}$ of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions such that $\mathrm{Y}$ is the set of their possible outcome states at $s$.
+The following fact, which is easy to verify, indicates that we can transform an action model to an actual/alpha neighborhood model without changing the satisfaction of formulas:
+Fact 1.
+Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
+  • Then $\mathtt{NM}=(\mathtt{ST},\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an actual neighborhood model such that for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
+  • Then $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{LE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an alpha neighborhood model such that for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
+Definition 8 (Action models $z$-representable by actual neighborhood models and $\alpha$-representable by alpha neighborhood models).
+  • An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $z$-representable by an actual neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
+    A class of action models $\mathbf{AM}$ is $z$-representable by a class of actual neighborhood models $\mathbf{NM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by an actual neighborhood model in $\mathbf{NM}$, and (2) every actual neighborhood model in $\mathbf{NM}$ $z$-represents an action model in $\mathbf{AM}$.
+  • An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $\alpha$-representable by an alpha neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{LE}_{\mathrm{C}}=\mathtt{nei}^{\alpha}_{\mathrm{C}}$.
+    A class of action models $\mathbf{AM}$ is $\alpha$-representable by a class of alpha neighborhood models $\mathbf{NM}^{\alpha}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}^{\alpha}$, and (2) every alpha neighborhood model in $\mathbf{NM}^{\alpha}$ $\alpha$-represents an action model in $\mathbf{AM}$.
 The following result can be easily shown:
@@
   1. (1)
-    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
-    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
+    For every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
   2. (2)
-    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}^{\alpha}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}^{\alpha}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
-    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}^{\alpha}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}^{\alpha}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}^{\alpha}$.
-3 Eight coalition logics determined by eight classes of grand-coalition-first action models
+    For every class of action models $\mathbf{AM}$ and class of alpha neighborhood models $\mathbf{NM}^{\alpha}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}^{\alpha}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}^{\alpha}$.
+Assume we have a coalition logic and the class of its models is $\mathbf{AM}$, a class of action models.
+By this result, if $\mathbf{AM}$ is representable by a class of neighborhood models $\mathbf{NM}$, then we can transform the action semantics of the coalition logic to the neighborhood semantics based on $\mathbf{NM}$.
+From Fact 1, from every action model in $\mathtt{AM}$, we can get a neighborhood model $\mathtt{NM}$ representing $\mathtt{AM}$.
+Let $\mathbf{NM}$ be the collection of all these neighborhood models.
+Clearly, $\mathbf{AM}$ is representable by $\mathbf{NM}$.
+Note that this transformation is not very useful.
+To make a transformation meaningful, $\mathbf{NM}$ shall be defined directly.
+In this paper, we mainly consider actual neighborhood models. In the sequel, by neighborhood models, we mean actual neighborhood models, unless explicitly stated.
+3 Eight coalition logics determined by eight classes of
+grand-coalition-first action models
 3.1 Eight classes of grand-coalition-first action models
 Definition 9 (Grand-coalition-first action models).
-A grand-coalition-first action model is a tuple $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, where:
-  • $\mathtt{ST}$ is a nonempty set of states.
-  • $\mathtt{AC}$ is a nonempty set of actions.
-  • $\mathsf{out}_{\mathtt{AG}}:\mathtt{ST}\times\mathtt{JA}_{\mathtt{AG}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathtt{AG}$.
-  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Definition 9 (Outcome functions and availability functions of grand-coalition-first action models).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
-For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
-$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}.$
-For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
-$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
-Definition 10 (Successor functions of grand-coalition-first action models).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},$ $\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
-Define the successor function $\mathtt{suc}$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ as follows: for all $s\in\mathtt{ST}$,
-$\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}.$
-For every set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$.
-A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$.
-It is easy to verify the following fact:
-Fact 2.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
-Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathtt{suc}(s)$.
-Fact 3.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
-Then:
-  1. (1)
-    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
-  2. (2)
-    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})$.
-Note that neither of the converses of the two statements holds. Figure 1 indicates a counterexample for them.
-$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
-Figure 1: This figure is used to show that in grand-coalition-first action models, the set of outcome states of a joint action might not contain the intersection of the sets of outcome states of the joint action’s parts.
-A grand-coalition-first action model is indicated in the figure, where we suppose $\mathtt{AG}=\{a,b\}$.
-It can be seen: (1) $\mathsf{out}_{a}(s_{0},\alpha_{1})$, the set of outcome states of $a$ performing $\alpha_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, (2) $\mathsf{out}_{a}(s_{0},\beta_{1})$, the set of outcome states of $b$ performing $\beta_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, but (3) $\mathsf{out}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$, the set of outcome states of $a$ and $b$ respectively performing $\alpha_{1}$ and $\beta_{1}$ at $s_{0}$, does not contain the intersection of $\mathsf{out}_{a}(s_{0},\alpha_{1})$ and $\mathsf{out}_{b}(s_{0},\beta_{1})$.
+An action model $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is a grand-coalition-first action model if for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
+It can be seen that for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}$ can be derived from $\mathsf{out}_{\mathtt{AG}}$, which is why we use “grand-coalition-first”.
 Definition 10 (Seriality, independence, and determinism of grand-coalition-first action models).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
 We say:
-  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
-  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
-  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, ${\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})}$ is a singleton.
+  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$ for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$;
+  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$ for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
+  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})$} is a singleton for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$.
 We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively, and let the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ correspond to the eight combinations of the three properties, respectively.
@@
 Tautologies ($\mathtt{A}\text{-}\mathtt{Tau}$): all propositional tautologies
-No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle[\mathrm{C}]\rangle\bot$
-Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{C}]\rangle\psi)$
-Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
+No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\bot$
+Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi)$
+Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
 Inference rules:
 Modus ponens ($\mathtt{MP}$): $\dfrac{\phi,\phi\rightarrow\psi}{\psi}$
-Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle[\mathrm{C}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi}$
-We let the following formulas respectively correspond to $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$.
-Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
-Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
-Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
+Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi}$
+We let the following formulas respectively correspond to $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$:
+Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\top$
+Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{C}\cup\mathrm{D}]\hskip-2.0pt\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
+Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle(\phi\lor\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi)$
 Definition 12 (Axiomatic systems for $\mathsf{MCL+X}$).
 For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inference rules of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
-Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [8]).
+Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [li_completeness_2024]).
 For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ given in Definition 12 is sound and complete with respect to the set of valid formulas of $\mathsf{MCL+X}$.
+3.3 Some facts about grand-coalition-first action models
+In this subsection, we present some facts about grand-coalition-first action models that will be used later.
+For every set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$.
+A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$.
+It is easy to verify the following:
+Fact 2.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
+Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
+Fact 3.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
+Then:
+  1. (1)
+    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$: $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
+  2. (2)
+    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{{a_{n}}})$.
+Neither of the converses of the two statements holds. What follows is a counterexample for both of them:
+Example 1.
+$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
+Figure 1: This figure is used to show that in grand-coalition-first action models, the set of outcome states of a joint action might not contain the intersection of the sets of outcome states of the joint action’s parts.
+We suppose $\mathtt{AG}=\{a,b\}$.
+A grand-coalition-first action model is indicated in Figure 1, where $(\alpha_{i},\beta_{j})$ represents the joint action of $a$ doing $\alpha_{i}$ and $b$ doing $\beta_{j}$.
+It can be seen: (1) $\mathsf{out}_{a}(s_{0},\alpha_{1})$, the set of outcome states of $a$ performing $\alpha_{1}$ at $s_{0}$, is equal to $\{s_{1},s_{2}\}$, (2) $\mathsf{out}_{b}(s_{0},\beta_{1})$, the set of outcome states of $b$ performing $\beta_{1}$ at $s_{0}$, is equal to $\{s_{1},s_{2}\}$, but (3) $\mathsf{out}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$, the set of outcome states of $a$ and $b$ respectively performing $\alpha_{1}$ and $\beta_{1}$ at $s_{0}$, does not contain the intersection of $\mathsf{out}_{a}(s_{0},\alpha_{1})$ and $\mathsf{out}_{b}(s_{0},\beta_{1})$.
 Remarks
-The class of grand-coalition-first $\mathsf{SID}$-models is $\alpha$-representable ([11, 5]), and the class of grand-coalition-first $\mathsf{SD}$-models is $\alpha$-representable ([14]).
-Are the other six classes of grand-coalition-first action models $\alpha$-representable?
-Are the eight classes of grand-coalition-first action models $z$-representable?
+The class of grand-coalition-first action $\mathsf{SID}$-models is $\alpha$-representable by a class of alpha neighborhood models meeting certain properties ([pauly_modal_2002, goranko_strategic_2013]).
+The class of grand-coalition-first action $\mathsf{SD}$-models is $\alpha$-representable by a class of alpha neighborhood models meeting certain properties ([shi_representation_2024]).
+Are the other six classes of grand-coalition-first action models $\alpha$-representable in a similar sense?
+Are the eight classes of grand-coalition-first action models $z$-representable in a similar sense?
 All this is yet unknown.
@@
 4.1 Outcome functions are compositional in many situations
-We say outcome functions of a class of action models are compositional if there is a way such that for every action model in the class, the outcome function of a coalition is determined by the outcome functions of its members in this way.
-Outcome functions in grand-coalition-first action models are not compositional. Figure 2 gives a counterexample for this.
+We say that outcome functions of a class of action models are compositional if there is a way such that for every action model in the class, the outcome function of a coalition is determined by the outcome functions of its members in this way.
+Outcome functions in grand-coalition-first action models are not compositional. What follows is a counterexample:
+Example 2.
+We suppose $\mathtt{AG}=\{a,b\}$.
+Two grand-coalition-first action models $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$ are depicted in Figure 2, where $\{\mathsf{out}^{1}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\{\mathsf{out}^{2}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$.
+In $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{1}\}$.
+In $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$: $\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{2}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{2}\}$.
+Note: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\mathsf{out}^{2}_{b}(s_{0},\beta_{1})$, but $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))\neq\mathsf{out}^{2}_{\mathtt{AG}}$ $(s_{0},(\alpha_{1},\beta_{1}))$.
+This implies that outcome functions in grand-coalition-first action models are not generally compositional: there is no way such that for every grand-coalition-first action model, the outcome function of a coalition is determined by the outcome functions of its members in this way.
 $s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{1}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
 $s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{2}$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
-Figure 2: This figure offers a counterexample that outcome functions in grand-coalition-first action models are not generally compositional.
-We suppose $\mathtt{AG}=\{a,b\}$.
-Two grand-coalition-first action models $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$ are depicted above, where $\{\mathsf{out}^{1}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\{\mathsf{out}^{2}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$.
-In $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{1}\}$ In $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$: $\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{2}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{2}\}$.
-Note: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\mathsf{out}^{2}_{b}(s_{0},\beta_{1})$, but $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))\neq\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$.
-This implies that outcome functions in grand-coalition-first action models are not generally compositional: there is no way such that for every grand-coalition-first action model, the outcome function of a coalition is determined by the outcome functions of its members in this way.
-There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other.
+Figure 2: This figure indicates two grand-coalition-first action models for Example 2.
+There are many situations in which every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other.
 In these situations, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
-What follows are two examples, and the second is from [3].
+What follows are two examples, and the second is from [alur_alternating-time_2002].
 Example 3.
 Adam and Bob control a ship lock with two doors: a front door and a back door.
-Adam can make the front door closed or open, Bob can make the back door closed or open, but they cannot make the two doors open at the same time.
+Adam can make the front door closed or open, Bob can make the back door closed or open, but they cannot make both doors open at the same time.
 This situation can be represented by the grand-coalition-first action model depicted in Figure 3, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
@@
 Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 3.
-Here, $c_{f}$ and $c_{b}$, respectively, express the front door is closed and the back door is closed.
+Here: $c_{f}$ and $c_{b}$, respectively, express the front door is closed and the back door is closed; $(\mathtt{open}\text{-}\mathtt{f},\mathtt{skip})$ represents a joint action where Adam closes the front door and Bob does nothing, and so on.
 It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
-For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{close}\text{-}\mathtt{f})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close}\text{-}\mathtt{f})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
+For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{close}\text{-}\mathtt{b})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close}\text{-}\mathtt{f})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
 Example 4.
 A system has two processes, $a$ and $b$.
-The process $a$ can assign $1$ or 00 to the variable $x$, and the process $b$ can assign $1$ or 00 to the variable $y$.
+The process $a$ can assign $1$ or 0 to the variable $x$, and the process $b$ can assign $1$ or 0 to the variable $y$.
 When $x=0$, $a$ can leave the value of $x$ unchanged or change it to $1$.
 When $x=1$, $a$ leaves the value of $x$ unchanged.
-Similarly, when $y=0$, $b$ can leave the value of $y$ unchanged or change it to $1$.
+Similarly, when $y=0$, $b$ can either leave the value of $y$ unchanged or change it to $1$.
 When $y=1$, $b$ leaves the value of $y$ unchanged.
@@
 For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\mathtt{AG}}(s_{4},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{4},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{4},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
-We want to mention that some literature, such as [6] and [7], deals with agency by propositional control: there are some agents; every atomic proposition is controlled by at most one agent.
+Remarks
+We want to mention that some literature, such as [harrenstein_boolean_2001] and [hoek_logic_2005], deals with agency by propositional control: there are some agents; every atomic proposition is controlled by at most one agent.
 In these settings, outcome functions are compositional.
@@
 Definition 13 (Single-coalition-first action models).
-A single-coalition-first action model is a tuple $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
-  • $\mathtt{ST}$ is a nonempty set of states.
-  • $\mathtt{AC}$ is a nonempty set of actions.
-  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
-  • for all $a\in\mathtt{AG}$, $\mathsf{out}_{a}:\mathtt{ST}\times\mathtt{JA}_{a}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $a$ such that for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
-  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Definition 15 (Outcome functions and availability functions of single-coalition-first action models).
-Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model.
-For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
-$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
-For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
-$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}$
+An action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is a single-coalition-first action model if the following conditions are met:
+  • for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$: $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
+  • for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{\mathrm{C}}}|_{a})\mid a\in\mathrm{C}\}$.
+It can be observed that for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}$ can be derived from $\mathsf{out}_{\emptyset}$ and all $\mathsf{out}_{a}$ where $a\in\mathrm{C}$.
+This is why we use “single-coalition-first”.
 Clearly, the outcome functions in single-coalition-first action models are compositional.
-For all $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, if for all $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}$ equals to $\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime}_{\mathrm{C}}|_{a}$.
-The following result offers two alternative definitions of sing-coalition-first action models.
+For all $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, if for all $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}$ is equal to $\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime}_{\mathrm{C}}|_{a}$.
+The following result offers two alternative definitions of single-coalition-first action models.
 Theorem 3.
-Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $s\in\mathtt{ST}$.
+Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $s\in\mathtt{ST}$.
 The following three sets of conditions are equivalent:
@@
     2. (b)
-      for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
-    It is easy to check this set of conditions is equivalent to the set of constraints on single-coalition-first action models given in Definition 14.
+      for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{\mathrm{C}}}|_{a})\mid a\in\mathrm{C}\}$.
   2. (2)
@@
     2. (b)
-      for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+      for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$: $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
   3. (3)
@@
 Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
-Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ $=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
+Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
 Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
 Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
-We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
+It suffices to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
 Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
@@
 By (1a) and (1b), it is easy to verify $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
-Assume ${\overline{\mathrm{C}}}=\emptyset$.
+Assume $\overline{\mathrm{C}}$$\emptyset$.
 Clearly, what we want holds.
@@
 By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$.
-Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$, which equals to $\mathsf{out}_{\emptyset}(s,\emptyset)$ by (3a).
+Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$, which is equal to $\mathsf{out}_{\emptyset}(s,\emptyset)$ by (3a).
 Let $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$.
@@
 Every single-coalition-first action model is a grand-coalition-first action model, but not vice versa.
-The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counterexample.
+The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model given in Example 1 is a counterexample.
 Remarks
-Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1].
-It can be verified convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
+Goranko and Jamroga [goranko_comparing_2004] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [abdou_rectangularity_1998].
+It can be verified that convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
 4.3 Single-coalition-first neighborhood models
@@
 Let $I$ be a nonempty set of indices.
-Define $\bigodot\{\Delta_{i}\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\}$ as expected.
+Define $\bigodot\{\Delta_{i}\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\}$ as expected [^3] Strictly speaking, we should write $\bigodot\Big\{(i,\Delta_{i})\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\Big\}$.
+Here, we abuse the notation a bit..
 Note $\bigodot\{\Delta\}=\Delta$.
 Definition 14 (Single-coalition-first neighborhood models).
-A single-coalition-first neighborhood model is a tuple $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
-  • $\mathtt{ST}$ is a nonempty set of states.
-  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
-  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$.
-  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
-Definition 17 (Neighborhood functions of single-coalition-first neighborhood models).
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
-For every $\mathrm{C}\subseteq\mathtt{AG}$, define the neighborhood function $\mathtt{nei}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
-$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}=\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}\neq\emptyset\end{cases}$
+A neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},$ $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is a single-coalition-first neighborhood model if the following conditions are met:
+  • for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\emptyset}(s)$ is either empty or a singleton with a nonempty element.
+  • for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
+  • for every nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$: $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
+The following fact will be used later.
 Fact 4.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
-Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
+Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
 Proof.
 Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
-Assume $\mathtt{suc}(s)=\emptyset$.
-Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is a cover of $\mathtt{suc}(s)$.
-Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$.
-Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is a cover of $\mathtt{suc}(s)$.
-Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$.
+Assume $\bigcup\mathtt{nei}_{\emptyset}(s)=\emptyset$.
+Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
+Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$.
+Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$, which is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
+Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$.
 Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
-Let $t\in\mathtt{suc}(s)$.
+Let $t\in\bigcup\mathtt{nei}_{\emptyset}(s)$.
 We want to show $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$.
-Note $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$ for every $a\in\mathtt{AG}$.
+Note $\mathtt{nei}_{a}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$ for every $a\in\mathtt{AG}$.
 Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in\mathrm{X}_{a_{1}}$, …, there is $\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in\mathrm{X}_{a_{n}}$.
@@
 Definition 15 (Seriality, independence, and determinism of single-coalition-first neighborhood models).
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
 We say:
   • $\mathtt{S}\textbf{-}\mathtt{NM}$ is serial if for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}(s)\neq\emptyset$.
-  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
+  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$: $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
   • $\mathtt{S}\textbf{-}\mathtt{NM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\mathrm{Y}\in\mathtt{nei}_{\mathtt{AG}}(s)$, $\mathrm{Y}$ is a singleton.
@@
 Remarks
-Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems.
-Goranko and Jamroga [4] discussed tight alternating transition systems, which we can verify are single-coalition-first neighborhood $\mathtt{SID}$-models.
+Alur, Henzinger, and Kupferman [alur_alternating-time_1998] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems.
+Goranko and Jamroga [goranko_comparing_2004] discussed tight alternating transition systems, which we can verify are single-coalition-first neighborhood $\mathtt{SID}$-models.
 4.4 Representation of single-coalition-first action models by single-coalition-first neighborhood models
@@
 Proof.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
-Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as follows:
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
+Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ as follows:
   • $\mathtt{AC}=\{\alpha_{a-s-\mathrm{X}}\mid a\in\mathtt{AG},s\in\mathtt{ST},\text{and }\mathrm{X}\in\mathtt{nei}_{a}(s)\}$.
+    • for every $s\in\mathtt{ST}$, $\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\mathtt{nei}_{\emptyset}(s)$.
     • for every $a\in\mathtt{AG}$, $s\in\mathtt{ST}$ and $\alpha_{x-u-\mathrm{X}}\in\mathtt{AC}$,
       $\mathsf{out}_{a}(s,\alpha_{x-u-\mathrm{X}})=\begin{cases}\mathrm{X}&\text{if $x=a$ and $u=s$}\\ \emptyset&\text{otherwise}\end{cases}$
-Let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
+    • for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{\mathrm{C}}}|_{a})\mid a\in\mathrm{C}\}$.
+It is easy to check that for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
+Then $\mathtt{S}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model.
 We claim that $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
 It suffices to show $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
-Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\text{ and }\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\}$.
-Assume $\mathtt{suc}(s)=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\emptyset$ and $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
-Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$.
-Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$ and $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$.
+Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\text{ and }\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\neq\emptyset\}$.
+Assume $\bigcup\mathtt{nei}_{\emptyset}(s)=\emptyset$.
+Then, $\mathsf{out}_{\emptyset}(s,\emptyset)=\emptyset$.
+Note for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
+Then, for all $a\in\mathtt{AG}$, $\sigma_{a}\in\mathtt{JA}_{a}$ and $s\in\mathtt{ST}$, $\mathsf{out}_{a}(s,\sigma_{a})=\emptyset$.
+Then, $\mathtt{AE}_{\mathrm{C}}(s)=\emptyset=\mathtt{nei}_{\mathrm{C}}(s)$.
+Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$.
+Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$ and $\mathtt{nei}_{\mathrm{C}}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$.
 Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
-Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
+Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$.
+Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
+Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
 Let $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$.
@@
 Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
-Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{a_{n}})$ and $\mathrm{X}\neq\emptyset$.
+Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$ and $\mathrm{X}\neq\emptyset$.
 Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\neq\emptyset$, …, $\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\neq\emptyset$.
@@
 Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$.
-Then, $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$.
-Let $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$.
+Then, $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s)$.
+Let $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s)$.
 Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$ for some $\mathrm{X}_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{n}\in\mathtt{nei}_{a_{n}}(s)$.
@@
 Proof.
-Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
 Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
 Definition 16 (Clear grand-coalition-first action models).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
-We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
-The following result offers two alternative definitions of clear grand-coalition-first action models.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
+We say that $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
+Intuitively, in clear grand-coalition-first action models, different action profiles have different outcome states.
+The following result indicates that in clear grand-coalition-first action models, different actions have different outcome states.
 Theorem 9.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model, and $s\in\mathtt{ST}$.
 The following three conditions are equivalent:
@@
 Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$, which is impossible.
-For any set $\mathrm{X}$ of states, we say $\Delta\subseteq\mathcal{P}(\mathrm{X})$ is a general partition of $\mathrm{X}$ if (1) $\bigcup\Delta=\mathrm{X}$ and (2) for every $\mathrm{Y},\mathrm{Y}^{\prime}\in\Delta$, if $\mathrm{Y}\neq\mathrm{Y}^{\prime}$, then $\mathrm{Y}\cap\mathrm{Y}^{\prime}=\emptyset$.
-Note a general partition $\Delta$ is a partition if $\emptyset\notin\Delta$.
+For any set $\mathrm{X}$ of states, we say that $\Delta\subseteq\mathcal{P}(\mathrm{X})$ is a general partition of $\mathrm{X}$ if (1) $\bigcup\Delta=\mathrm{X}$, and (2) for every $\mathrm{Y},\mathrm{Y}^{\prime}\in\Delta$, if $\mathrm{Y}\neq\mathrm{Y}^{\prime}$, then $\mathrm{Y}\cap\mathrm{Y}^{\prime}=\emptyset$.
+A general partition $\Delta$ is a partition if $\emptyset\notin\Delta$.
 Fact 5.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
-Then, if $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, then for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
+Then, if $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, then for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
 Proof.
 Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear.
-Let $\mathrm{C}\subseteq\mathtt{AG}$ and $t\in\mathtt{ST}$.
-Note $\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$.
-It is easy to show $\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}=\mathtt{suc}(s)$.
+Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
+Note $\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$.
+It is easy to show $\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$.
 Let $\mathrm{X}_{1},\mathrm{X}_{2}\in\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ such that $\mathrm{X}_{1}\neq\mathrm{X}_{2}$.
@@
 Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$, that is, $\mathrm{X}_{1}\cap\mathrm{X}_{2}=\emptyset$.
-Note the other direction of the statement of this fact might not hold.
+Note that the other direction of the statement of this fact might not hold. What follows is a counterexample.
+Example 5.
 Suppose $\mathtt{AG}=\{a\}$.
-Consider a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ such that:
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model such that:
   • $\mathtt{ST}=\{s\}$;
+  • $\mathsf{out}_{\emptyset}(s,\emptyset)=\mathtt{ST}$;
   • $\mathtt{AC}=\{\alpha_{1},\alpha_{2}\}$;
   • $\mathsf{out}_{a}(s,\alpha_{1})=\mathsf{out}_{a}(s,\alpha_{2})=\mathtt{ST}$.
-It is easy to check for all $\mathrm{C}\subseteq\mathtt{AG}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$.
+It is easy to check for all $\mathrm{C}\subseteq\mathtt{AG}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
 However, $\mathtt{G}\textbf{-}\mathtt{AM}$ is not clear.
@@
 Proof.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a clear grand-coalition-first action model.
 To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model, we need to show:
   1. (1)
-    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
+    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
   2. (2)
-    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
-    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
-  3. (3)
-    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$,
-    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
-By Fact 2, (1) holds. Clearly, (3) holds. It remains to show (2). Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
-Assume $\mathrm{C}=\emptyset$.
-Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{\mathtt{AG}}\}=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}=\mathtt{suc}(s)$;
-Assume $\mathrm{C}\neq\emptyset$.
+    for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
+    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$
+By Fact 2, (1) holds. It remains to show (2). Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. Assume $\mathrm{C}\neq\emptyset$
 Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
@@
 Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
-It is easy to find a single-coalition-first action model that is not clear.
+It is easy to see that the action model given in Example 5 is a single-coalition-first action model that is not clear.
 Remarks
-Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models.
-As observed by Ågotnes and Alechina [12], the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
+Goranko and Jamroga [goranko_comparing_2004] and Ågotnes and Alechina [agotnes_embedding_2015] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models.
+As observed by Ågotnes and Alechina [agotnes_embedding_2015], the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
+The reason that Goranko and Jamroga [goranko_comparing_2004] and Ågotnes and Alechina [agotnes_embedding_2015] used “injective” is as follows: in injective concurrent game models, the outcome function $\mathsf{out}_{\mathtt{AG}}$ for $\mathtt{AG}$ is an injective function with the set of possible states as its range.
+However, in grand-coalition-first models, the outcome function $\mathsf{out}_{\mathtt{AG}}$ for $\mathtt{AG}$ takes the power set of the set of possible states as its range.
+A grand-coalition-first action model where the outcome function $\mathsf{out}_{\mathtt{AG}}$ for $\mathtt{AG}$ is injective does not imply that it is clear.
+This is why we do not use “injective” for clear grand-coalition-first action models.
 5.2 Clear single-coalition-first neighborhood models
 Definition 17 (Clear single-coalition-first neighborhood models).
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
-We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
+We say that $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
 The following result gives an alternative definition of clear single-coalition-first neighborhood models.
 Theorem 11.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
 The following two conditions are equivalent:
   1. (1)
-    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
+    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$;
   2. (2)
-    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
+    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
 Proof.
@@
 Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
-By Fact 4, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$.
-Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$.
+By Fact 4, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
+Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
 Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
@@
 Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$.
-It is easy to see $\mathtt{suc}(s)\neq\emptyset$.
-Then, $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$, which is a singleton.
+It is easy to see $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$.
+Then, $\mathtt{nei}_{\emptyset}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$, which is a singleton.
 Note $\mathtt{nei}_{\mathrm{C}}(s)$ is not a singleton.
@@
 By Theorem 9, clear grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition.
-It might be expected that the following condition, which concerns the grand coalition, is equivalent to the two conditions given in Theorem 11: for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$.
+It might be expected that the following condition, which concerns the grand coalition, is equivalent to the two conditions given in Theorem 11: for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
 Actually, this is not the case.
 What follows is a counterexample.
+Example 6.
 Assume $\mathtt{AG}=\{a,b\}$.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model such that:
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model such that:
   • $\mathtt{ST}=\{s_{0},s_{1},s_{2},s_{3}\}$;
-  • $\mathtt{suc}(s_{0})=\{s_{1},s_{2},s_{3}\}$;
+  • $\mathtt{nei}_{\emptyset}(s_{0})=\{\{s_{1},s_{2},s_{3}\}\}$;
   • $\mathtt{nei}_{a}(s_{0})=\{\{s_{1},s_{2}\},\{s_{2},s_{3}\}\}$ and $\mathtt{nei}_{b}(s_{0})=\{\{s_{2}\},\{s_{1},s_{3}\}\}$.
-Note $\mathtt{nei}_{a}(s_{0})$ is not a partition of $\mathtt{suc}(s_{0})$.
-However, it can be checked that $\mathtt{nei}_{\mathtt{AG}}(s_{0})=\{\{s_{1}\},\{s_{2}\},\{s_{3}\}\}$, which is a partition of $\mathtt{suc}(s_{0})$.
+Note $\mathtt{nei}_{a}(s_{0})$ is not a partition of $\bigcup\mathtt{nei}_{\emptyset}(s_{0})$.
+However, it can be checked that $\mathtt{nei}_{\mathtt{AG}}(s_{0})=\{\{s_{1}\},\{s_{2}\},\{s_{3}\}\}$, which is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s_{0})$.
 Remarks
-Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems.
+Alur, Henzinger, and Kupferman [alur_alternating-time_1998] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems.
 It can be shown that they are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
@@
 It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model.
-∎
 Theorem 13.
@@
 Proof.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model.
-Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model.
+Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6.
 As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
@@
 Definition 18 (Histories in grand-coalition-first action models).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
 For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$, where every $s_{i}$ is a state and every $\sigma_{\mathrm{C}}^{i}$ is a joint action of $\mathrm{C}$, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\sigma_{\mathrm{C}}^{i+1}\in\mathtt{av}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathsf{out}_{\mathrm{C}}(s_{i},\sigma_{\mathrm{C}}^{i+1})$.
-Specially, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
+Specially, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $\in\mathtt{ST}, ss$ is called a $\mathrm{C}$-history from $s$ to $s$.
 Definition 19 (Tree-like grand-coalition-first action models).
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
-We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
-Note every tree-like model has a unique root.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
+We say that $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
+Note that every tree-like model has a unique root.
 The following result gives two alternative definitions of tree-like grand-coalition-first action models.
 Theorem 15.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
 The following three conditions are equivalent:
   1. (1)
-    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
+    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$;
   2. (2)
-    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
+    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $\mathrm{C}$-history from $r$ to $s$;
   3. (3)
-    there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
+    there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
 Proof.
@@
 Assume (1).
-Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$.
+Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
 Let $s\in\mathtt{ST}$.
-We want to show that there is an unique $\mathtt{AG}$-history from $r$ to $s$.
+We want to show that there is a unique $\mathtt{AG}$-history from $r$ to $s$.
 Let $a\in\mathtt{AG}$.
@@
 Let $\theta_{a}$ be the unique $a$-history from $r$ to $s$.
-Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
+Note that it is impossible that $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
 Assume $\theta_{a}=(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$.
@@
 Assume there is a different $\mathtt{AG}$-history $\theta^{\prime}_{\mathtt{AG}}$ from $r$ to $s$.
-Note it is impossible $\theta^{\prime}_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$.
+Note that it is impossible that $\theta^{\prime}_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$.
 Let $\theta^{\prime}_{\mathtt{AG}}=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
@@
 Assume (3).
-Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
+Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
 Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
-We want to show that there is an unique $\mathrm{C}$-history from $r$ to $s$.
+We want to show that there is a unique $\mathrm{C}$-history from $r$ to $s$.
 Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$.
 Assume there is a $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$ which is different from $r$.
-Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Note that it is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
 Assume $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$.
@@
 Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$.
-Note it is impossible $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$.
+Note that it is impossible that $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$.
 Let $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$.
@@
 Assume there is a different $\mathrm{C}$-history $\theta^{\prime}_{\mathrm{C}}$ from $r$ to $s$.
-Again, it is impossible $\theta_{\mathrm{C}}^{\prime}=x$ for some $x\in\mathtt{ST}$.
+Again, it is impossible that $\theta_{\mathrm{C}}^{\prime}=x$ for some $x\in\mathtt{ST}$.
 Let $\theta^{\prime}_{\mathrm{C}}=(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$.
-For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$.
+For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\lambda^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$.
 Then, $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
@@
 Proof.
-Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model and $r$ be its root.
+Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a tree-like grand-coalition-first action model and $r$ be its root.
 Let $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$.
@@
 Remarks
-Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models, which are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined in this section.
+Ågotnes, Goranko, and Jamroga [agotnes_alternating-time_2007] defined tree-like concurrent game models, which are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined in this section.
 For the former, it is only required that for every state, there is a unique computation, which is a sequence of states, from the root to the state.
@@
 Definition 20 (Histories in single-coalition-first neighborhood models).
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
 For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
-Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
+Specifically, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $\in\mathtt{ST}, ss$ is called a $\mathrm{C}$-history from $s$ to $s$.
 Definition 21 (Tree-like single-coalition-first neighborhood models).
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
-We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
-The following result is easy to verify and we skip its proof.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
+We say that $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
+The following result is easy to verify, and we skip its proof.
 Fact 6.
@@
 Theorem 17.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
 The following two conditions are equivalent:
   1. (1)
-    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
+    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$;
   2. (2)
-    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
+    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $\mathrm{C}$-history from $r$ to $s$.
 Proof.
@@
 Assume (1).
-Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
+Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and every $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
 Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
@@
 Assume there is a different $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$.
-Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Note that it is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
 Let $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
 Assume $\mathrm{C}=\emptyset$.
-Then for every $i$ such that $1\leq i\leq n$, $\mathrm{Y}_{i}=\mathtt{suc}(s_{i-1})$.
+Then for every $i$ such that $1\leq i\leq n$, $\mathrm{Y}_{i}=\bigcup\mathtt{nei}_{\emptyset}(s_{i-1})$.
 Let $a\in\mathtt{AG}$.
-For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$.
+For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$.
 Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$.
@@
 Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$ for every $s\in\mathtt{ST}$.
-For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq\mathrm{Z}_{i}$.
-Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{m},s_{m})$ is an $a$-history from $r$ to $r$, which is different from $r$.
+For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq\mathrm{Z}_{i}$.
+Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$.
 Note $r$ is an $a$-history from $r$ to $r$.
@@
 Let $a\in\mathtt{AG}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$.
-It is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
+It is impossible that $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
 Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
-It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),$ $s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
+It is easy to see $(s_{0},\bigcup\mathtt{nei}_{\emptyset}(s_{0}),s_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(s_{n-1}),$ $s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
 Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$.
-We want to show $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
-It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
-Let $\theta_{\mathrm{C}}=(t_{0},\mathtt{suc}(t_{0}),t_{1},\dots,\mathtt{suc}(t_{m-1}),t_{m})$.
-For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathtt{suc}(t_{i-1})$ and $t_{i}\in\mathrm{Z}_{i}$.
+We want to show $\theta_{\mathrm{C}}=(s_{0},\bigcup\mathtt{nei}_{\emptyset}(s_{0}),s_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(s_{n-1}),s_{n})$.
+It is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Let $\theta_{\mathrm{C}}=(t_{0},\bigcup\mathtt{nei}_{\emptyset}(t_{0}),t_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(t_{m-1}),$ $t_{m})$.
+For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\bigcup\mathtt{nei}_{\emptyset}(t_{i-1})$ and $t_{i}\in\mathrm{Z}_{i}$.
 Then, $(t_{0},\mathrm{Z}_{1},t_{1},\dots,\mathrm{Z}_{m},t_{m})$ is an $a$-history from $r$ to $s$.
 Then $n=m$ and $s_{i}=t_{i}$ for all $0\leq i\leq n$.
-Then, $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
+Then, $\theta_{\mathrm{C}}=(s_{0},\bigcup\mathtt{nei}_{\emptyset}(s_{0}),s_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(s_{n-1}),s_{n})$.
 Assume $\mathrm{C}\neq\emptyset$.
 Let $a\in\mathrm{C}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$.
-Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
+Note that it is impossible that $\theta_{a}=x$ for some $x\in\mathtt{ST}$.
 Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
 For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$.
-Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
-Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$.
-It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
+Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
+Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$.
+It is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$.
 Let $\theta_{\mathrm{C}}=(t_{0},\mathrm{U}_{1},t_{1},\dots,\mathrm{U}_{m},t_{m})$.
@@
 In fact, this is not the case.
-The example given after Theorem 11 also works here.
-6.3 Representation of tree-like grand-coalition-first action models by tee-like single-coalition-first neighborhood models
+Example 6 also works here.
+6.3 Representation of tree-like grand-coalition-first action models by tree-like single-coalition-first neighborhood models
 Theorem 18.
@@
 It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model.
-∎
 Theorem 19.
@@
 Proof.
-Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model.
-Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6.
+Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model.
+Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6.
 As shown there, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
@@
 7 Each of those eight coalition logics is determined by these six kinds of models, too
-Theorem 21.
-For every $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, there is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime})$ such that for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime}\Vdash\phi$.
-Proof.
-Lert $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
 In the following, for every finite nonempty sequence $\pi$, we use $\pi^{l}$ to denote the last element of $\pi$.
-Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},$ $\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ as follows:
+Definition 22 (Unravelling of pointed grand-coalition-first action models).
+Let $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action model, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
+Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label}^{\prime})$ as follows:
   • $\mathtt{ST}^{\prime}=\mathtt{ST}_{0}^{\prime}\cup\mathtt{ST}_{1}^{\prime}\cup\mathtt{ST}_{2}^{\prime}\cup\dots$, where:
@@
     • for every $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})=\{\pi-\sigma_{\mathtt{AG}}-u\mid u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\}$.
+    • for all $\mathrm{C}\subseteq\mathtt{AG}$, $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
+  • for every $\mathrm{C}\subseteq\mathtt{AG}$ and $\pi\in\mathtt{ST}^{\prime}$, $\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{{\mathrm{C}}})\neq\emptyset\}$.
   • for every $\pi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\pi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\pi^{l})\}$.
-Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$ and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
-We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is what we want.
-We first make two claims:
+$(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is called the unravelling of $(\mathtt{G}\textbf{-}\mathtt{AM},s)$.
+It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is tree-like.
+This notion is a variation of the notion of unravelling used throughout modal logic.
+The slight difference is that, in the unravelled pointed model defined above, states are sequences consisting of states (of the original model) and action profiles, rather than just states (of the original model).
+Lemma 1.
+Let $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action model and $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ be its unravelling, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ and $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
+Then:
   1. (1)
@@
     for every $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, then there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
-First, we show the claim (1).
+Proof.
+  1. (1)
     Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$.
@@
 Then, $u\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, that is, $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
-Second, we show the claim (2).
+  2. (2)
     Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $t\in\mathtt{ST}$.
@@
 Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
-We show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is an $\mathtt{X}$-model.
-[…]
+Lemma 2.
+Let $\mathtt{X}\in\mathrm{ES}$, $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model, and $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ be the unravelling, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ and $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
+Then $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is a pointed grand-coalition-first action $\mathtt{X}$-model.
+Proof.
+Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial.
+Let $\pi\in\mathtt{ST}^{\prime}$.
+Note there is $\sigma_{\mathtt{AG}}$ such that $\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$ $\neq\emptyset$.
+Let $u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$.
+Then, $\pi-\sigma_{\mathtt{AG}}-u\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$.
+Then, $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})\neq\emptyset$.
+Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is serial.
 Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent.
@@
 Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic.
-Then, there is $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\pi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ is not a singleton.
+Then, there is $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}^{\prime}_{\mathtt{AG}}(\pi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ is not a singleton.
 Let $\pi^{\prime},\pi^{\prime\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ such that $\pi^{\prime}\neq\pi^{\prime\prime}$.
@@
 We have a contradiction.
-It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is tree-like.
-We show for all $\pi\in\mathtt{ST}^{\prime}$ and $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\phi$, which implies for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$.
-We put an induction on $\phi$.
-We consider only the case $\phi=\langle[\mathrm{C}]\rangle\psi$ and skip others.
-Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\langle[\mathrm{C}]\rangle\psi$.
+Lemma 3.
+Let $\mathtt{X}\in\mathrm{ES}$, $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action model, and $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ be the unravelling of $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ and $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
+Then for all $\pi\in\mathtt{ST}^{\prime}$ and $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\phi$.
+Proof.
+We put an induction on $\phi$. We consider only the case $\phi=\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi$ and skip others.
+Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi$.
 Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ such that for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
-We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
+We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
 Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$.
@@
 By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
-Assume $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\langle[\mathrm{C}]\rangle\psi$.
+Assume $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi$.
 Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
@@
 By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
+From Lemmas 2 and 3, the following result follows:
+Theorem 21.
+For every $\mathtt{X}\in\mathrm{ES}$ and pointed grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, there is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime})$ such that for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime}\Vdash\phi$.
 Theorem 22.
@@
 By Theorem 20 and Theorem 1, (5) and (6) are equivalent.
-It suffices to show that (0), (1), (3) and (5) are equivalent.
+It suffices to show that (0), (1), (3), and (5) are equivalent.
 By Theorem 4, (0) implies (1).
@@
 8 Conclusion
-Li and Ju [9, 8] proposed eight coalition logics based on grand-coalition-first action models.
-In this work, we show that each of them is determined by six kinds of models, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
+Li and Ju [li_minimal_2025, li_completeness_2024] proposed eight coalition logics based on grand-coalition-first action models.
+In this work, we show that each of them is also determined by six other kinds of models: single-coalition-first action models, single-coalition-first actual neighborhood models, clear grand-coalition-first action models, clear single-coalition-first actual neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first actual neighborhood models.
+Among the six kinds of models, except for tree-like grand-coalition-first action models and tree-like single-coalition-first actual neighborhood models, which mostly make technical sense, the other four kinds of models make good sense for representing coalitional powers.
+The language of coalition logic is not very expressive and cannot distinguish between them.
+Further work worth doing is to consider richer languages.
+The operator $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ of coalition logic is a bundled operator: it implicitly has an existential quantifier over actions, a universal operator over possible outcomes of actions, and a next-time operator.
+What follows are three ideas.
+The first is to consider the language of Alternating-time Temporal Logic $\mathsf{ATL}$ and to see whether we could have similar results.
+The language of $\mathsf{ATL}$ is an extension of the language of coalition logic with the operators $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\mathtt{G}\phi$ ($\mathrm{C}$ has a strategy to ensure $\phi$ forever) and $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\mathtt{U}\psi$ ($\mathrm{C}$ has a strategy to ensure $\phi$ until $\psi$).
+The second idea is to separate the quantifier over actions in the meaning of $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$.
+Logics following this idea would have connections to logics of “bring-it-about” [porn_logic_1970, elgesem_action_1993].
+The third idea is to consider languages of STIT (see to it that) logics [belnap_seeing_1988], which are temporal extensions of logics of “bring-it-about”.
 Acknowledgments
-Thanks go to Thomas Ågotnes, Valentin Goranko, and the audience of a logic seminar at Beijing Normal University.
+Thanks go to Thomas Ågotnes, Valentin Goranko, the three anonymous reviewers and the audience of the 6th International Conference on Logic and Argumentation, and the audience of a logic seminar at Beijing Normal University.